### v1

# Pariksha : A Large-Scale Investigation of Human-LLM Evaluator Agreement on Multilingual and Multi-Cultural Data

Abstract
Evaluation of multilingual Large Language Models (LLMs) is challenging due to a variety of factors -- the lack of benchmarks with sufficient linguistic diversity, contamination of popular benchmarks into LLM pre-training data and the lack of local, cultural nuances in translated benchmarks. In this work, we study human and LLM-based evaluation in a multilingual, multi-cultural setting. We evaluate 30 models across 10 Indic languages by conducting 90K human evaluations and 30K LLM-based evaluations and find that models such as GPT-4o and Llama-3 70B consistently perform best for most Indic languages. We build leaderboards for two evaluation settings - pairwise comparison and direct assessment and analyse the agreement between humans and LLMs. We find that humans and LLMs agree fairly well in the pairwise setting but the agreement drops for direct assessment evaluation especially for languages such as Bengali and Odia. We also check for various biases in human and LLM-based evaluation and find evidence of self-bias in the GPT-based evaluator. Our work presents a significant step towards scaling up multilingual evaluation of LLMs. [^1] We plan to release all evaluation artifacts post acceptance
[…]
[backgroundcolor=gray!10, linecolor=black, outerlinewidth=0.5pt, roundcorner=1mm, skipabove=skipbelow=font=, ]promptbox
Pariksha : A Large-Scale Investigation of Human-LLM Evaluator Agreement on Multilingual and Multi-Cultural Data
[…]
wattsishaan18@gmail.com, sunayana.sitaram@microsoft.com
1 Introduction
Figure 1: Evaluation pipeline (1) We curate a diverse set of evaluation prompts with the help of native speakers. (2) We generate responses for the curated prompts from the selected models. (3) We evaluate generated responses in two settings (direct assessment and pairwise comparison) by both Humans and an LLM. (4) We construct leaderboards using scores obtained and analyze the agreement between human and LLM evaluators.
Large Language Models (LLMs) have made tremendous progress recently by excelling at several tasks (OpenAI et al., 2024; Zhang et al., 2023; Anil and Team, 2024; Reid and Team, 2024, interalia). However, it is not always clear what capabilities these models possess, leading to an increased interest in evaluation. Benchmarking is the defacto standard for evaluating LLMs, with several popular benchmarks used to validate the quality of models when they are released.
However, standard benchmarking suffers from the following issues: many popular benchmarks are available on the web and have already been consumed in the training data of LLMs, rendering them unsuitable for fair evaluation. This phenomenon is known as test dataset contamination, and recent work Ravaut et al. (2024); Golchin and Surdeanu (2024); Dong et al. (2024); Oren et al. (2024); Deng et al. (2024) has suggested that contamination can occur not only during pre-training, but also during fine-tuning and evaluation Balloccu et al. (2024). This calls for dynamic benchmarking with the help of humans Chiang et al. (2024); Yang et al. (2023). Although, human evaluation is considered the gold standard, it can be expensive and time consuming. Due to this, the use of LLM-evaluators, where an LLM itself is used to evaluate the output of another LLM (sometimes the same LLM) has become very popular.
Most studies on LLM training and evaluation focus on English. Recent work has shown that LLMs perform worse on non-English languages, particularly those written in scripts other than the Latin script, and under-resourced languages Ahuja et al. (2023, 2024); Asai et al. (2024). Studies on cultural values in LLMs have also shown that frontier models such as GPT-4 align more closely to Western, Rich, Industrialized norms Rao et al. (2023). This has led to a proliferation of models being built for specific languages, cultures and regions such as Indic, Arabic, African, Chinese, European, and Indonesian (Gala et al., 2024; Sengupta et al., 2023; Zeng et al., 2023; Bai et al., 2023; Jiang et al., 2023, 2024; Cahyawijaya et al., 2024; Cohere, 2024, interalia). Multilingual evaluation is challenging due to the small number of multilingual benchmarks available, the lack of language diversity in them Ahuja et al. (2022) and the evidence of possible contamination of many of these benchmarks Ahuja et al. (2024). Additionally, many multilingual benchmarks are translations of benchmarks originally created in English, leading to loss of linguistic and cultural context.
In this work, we perform 90K human evaluations - the largest scale multilingual human evaluation of LLMs as per our knowledge. We perform evaluation on a new set of general and culturally nuanced prompts created independently by native speakers for each language. We use a setting similar to the LMSys ChatbotArena Chiang et al. (2024) and ask human evaluators employed by Karya to perform two evaluation tasks: comparative evaluations between models, and individual evaluations or direct assessments of 30 models. Karya employs workers from all states of India, with a focus on rural and marginalized communities, making our study the first effort as per our knowledge that includes these communities in the evaluation process. In addition to performing human evaluations, we build upon prior work on LLMs as multilingual evaluators Hada et al. (2024b, a) to perform the same evaluations using LLMs as judges. We also use LLMs to perform safety evaluation, for which we do not engage Karya workers due to ethical concerns.
Our contributions are as follows: (1) We perform 90K human evaluations across 10 Indic languages, comparing 30 Indic and multilingual models using pairwise and direct assessment on a culturally-nuanced dataset. (2) We perform the same evaluations using an LLM-based evaluator to analyze the agreement between human and LLM evaluation, making this work the most comprehensive analysis of LLM-based evaluators in the multilingual setting. (3) We create leaderboards based on human and LLM-based evaluators and analyze trends and biases across languages and models.
2 Related Work
Multilingual Evaluation Benchmarks
Ahuja et al. (2023, 2024); Asai et al. (2024) conduct comprehensive multilingual evaluations of open-source and proprietary models on a large scale across various available multilingual benchmarks. Liu et al. (2024) release a Multilingual Generative test set that can assess the capability of LLMs in five different languages. Other popular multilingual NLU benchmarks include XGLUE Liang et al. (2020), XTREME Hu et al. (2020), XTREME-R Ruder et al. (2021).
Indic Evaluation Benchmarks
Kakwani et al. (2020) release the first Indic NLU benchmark, IndicGLUE, for 11 languages. Doddapaneni et al. (2023) build on top of the former and release IndicXTREME, spanning all 22 languages. On the NLG side, Kumar et al. (2022) offer IndicNLGsuite, covering 5 tasks across 11 languages. Gala et al. (2023) release a machine translation benchmark, IN22, for both conversational and general translation evaluation across all 22 languages. Recently, Singh et al. (2024a) put forth IndicNLGBench, a collection of diverse generation tasks like cross-lingual summarization, machine translation, and cross-lingual question answering.
Human Evaluation
Several previous studies have used humans to evaluate LLMs, build leaderboards, or as strong upper-bound baselines Chiang et al. (2024); Wu and Aji (2023); Srivastava and Team (2023); Hada et al. (2024b, a); Chiang and Lee (2023). Others have employed humans to create gold-standard culturally-nuanced evaluation prompts or to evaluate the corresponding outputs of various LLMs Singh et al. (2024b); Üstün et al. (2024); Cahyawijaya et al. (2024); Feng et al. (2024)
LLM-based Automatic Evaluations
LLMs have been shown to be useful as evaluators due to their instruction following abilities, but studies have also shown that they can be biased and may not always agree with human judgments. Hada et al. (2024b, a) conduct a comprehensive survey of LLMs as an evaluators in the multilingual setting, and also release, METAL, a benchmark for LLM-based Summarization evaluation across 10 languages. Other recent works such as Liu et al. (2024); Shen et al. (2023); Kocmi and Federmann (2023) also discuss and use LLMs for evaluations at scale, and Zheng et al. (2023) employ GPT-4 as an evaluator alongside humans to build the MTBench and ChatbotArena leaderboard. Ning et al. (2024) propose an LLM-based peer-review process to automatically evaluate the outputs of an LLM, by other models in the setup.
3 Methodology
Our evaluation setup is summarized in Figure 1.
3.1 Prompt Curation
We include the following 10 Indian languages in our evaluation: Hindi, Tamil, Telugu, Malayalam, Kannada, Marathi, Odia, Bengali, Gujarati, and Punjabi. Our prompts comprise of 20 questions per language - 5 on health, 5 on finance, and 10 culturally nuanced prompts that were created independently by native speakers in the research team and Karya . Although we currently evaluate only on a small set of prompts, we plan to scale the number of prompts by allowing evaluators to create their own prompts similar to ChatbotArena Yang et al. (2023).
3.2 Model Selection
We evaluate popular Indic language models in addition to the leading proprietary LLMs. Most of the Indic LLMs are fine-tuned versions of the open-source Llama-2 7B base model Touvron et al. (2023), Mistral 7B Jiang et al. (2023) or Gemma 7B Mesnard and Team (2024) models, hence we added the instruct versions of these models to our evaluation to determine the gain obtained by fine-tuning these models with Indic data. We have also included the latest Llama-3 8B and Llama-3 70B AI@Meta (2024) models to evaluate their effectiveness for multilingual fine-tuning. We list all models under consideration Appendix B in Table 2 and Table 3. We are aware that it is not entirely fair to compare open-source models with API-based systems that may have several other components in place, such as language detectors, more sophisticated safety guardrails etc., however, we treat all models as the same for this study and urge the reader to keep this in mind while interpreting the results. The details for generating model query-response pairs can be found in Appendix C.
3.3 Evaluation Setup
We evaluate the generated model query-response using two different strategies and by two types of evaluators. First, we do a pairwise comparison (battle) between model responses for the same prompt and calculate Elo Ratings Elo (1978); Boubdir et al. (2023). Second, we also calculate various direct assessment metrics for each model prompt-response data point. We evaluate 12-15 models for each language except Hindi for which we evaluate 20 models. The detailed statistics of the evaluation datapoints can be seen in Appendix F.
We use human evaluation to annotate the datapoints. Each datapoint is annotated by three annotators and the majority vote is taken. If all three votes are different, we treat it as a tie in case of a battle and take average score in case of direct assessment metrics. We also use an LLM (GPT-4-32K) for annotating the battles as well as calculating the direct assessment metrics. The instructions are provided in English and a detailed description of the task and scoring rubric is also provided.
3.3.1 Pairwise comparison
We use the Elo Rating systems, which is widely used in chess to measure the relative skills of players. This helps us to convert human preferences into scores, which can predict the win rates between different models. This system is also employed in the LMSys Chatbot Arena setup [^2] https://huggingface.co/spaces/lmsys/chatbot-arena-leaderboard Chiang et al. (2024). Additionally, we employ the Maximum Likelihood Estimation (MLE) Elo rating system to determine rankings, as it remains unaffected by the sequence of comparisons. More information about Elo is available in Appendix A.
Battle Generation
We generate ($N\choose 2$$\times$ (number of prompts) pairwise comparisons for each language. To check for annotator and LLM consistency, we added duplicate pairings with responses flipped for 10% of the original pairings. The battles were designed in such a way that each model contributed to Response A and Response B equally. The detailed statistics of datapoints can be seen in Table 4. For pairwise comparisons, we evaluate 21690 datapoints using three human annotators and the LLM-evaluator.
Human evaluation setup
The annotators perform the evaluation task on a smartphone. The annotators are provided with the query, the two model responses (model names are hidden), and set of three options - A (response 1 is better), B (response 2 is better), and C (tie, equally good/bad). We also ask the annotators to provide a spoken justification for the chosen response that is captured as audio by the app. The annotation guidelines and Hindi app screenshots are available in Appendix E.1.
LLM evaluation setup
We also evaluate battles using GPT-4-32K as an LLM evaluator. The setting is similar to the one provided to humans. The detailed prompt is provided in Figure 10.
3.3.2 Direct Assessment
In addition to a pairwise comparison, humans as well as the LLM also rate a query-response pair on three metrics - Linguistic Acceptability (LA), Task Quality (TQ), and Hallucination (H) metrics Hada et al. (2024b, a). We evaluate a total of 8640 datapoints across the 3 metrics and the 10 languages, detailed statistics can be seen in Table 4. We rank each model based on the average scores obtained across all query-response pairs with 5 being the maximum (2LA + 2TQ + 1H) and 0 being the lowest possible score.
Human evaluation setup
The annotators are shown the query-response pair and a checkbox asking if the output is gibberish. If selected the response is automatically given the lowest score, otherwise, the annotators are asked to label the three metrics. The annotation guidelines and Hindi app screenshots are available in Appendix E.2.
LLM evaluation setup
For LLM-based evaluation, we make a single call for each metric using the prompt in Fig 11 resulting in a total of 3 calls per model per query. The detailed description for each metric rubric can be found in Figures 12, 13 and 14. Our metric prompts were sourced from Chiang et al. (2024); Hada et al. (2024b, a) and tailored to our use-case.
3.3.3 Safety Evaluation
We use the Hindi prompts from RTP-LX de Wynter et al. (2024) dataset [^3] https://github.com/microsoft/RTP-LX which is specifically designed to elicit toxic responses and ask the relevant models to generate completions. These completions are then evaluated using an LLM evaluator with the same prompt used for individual evaluations (Fig 11). The detailed rubric for Safety is defined in Fig 15. We also perform an exact match with the Hindi block words from the FLORES Toxicity-200 dataset [^4] https://github.com/facebookresearch/flores/blob/main/toxicity/README.md Costa-jussà et al. (2022) to check for toxic words in the output.
3.4 Inter-Annotator Agreement
To check for the quality of human annotation, we calculate inter-annotator agreement between the three human annotators using two metrics - Percentage Agreement (PA) and Fleiss Kappa ($\kappa$). These metrics are also used to judge the alignment between humans and LLMs for the evaluation tasks, following the same setup as prior work Hada et al. (2024b, a). We also calculate the correlation between rankings of the leaderboards obtained from human and LLM evaluations using Kendall’s Tau ($\tau$).
4 Results
4.1 Leaderboard Analysis
Leaderboard Setup
Figure 2 depicts a visualization of the leaderboard based on the MLE Elo rating method discussed in Section 3.3.1. For the Direct Assessment scores, we report the average score across all query-response pairs for a model in Figure 3. We include both human and LLM-evaluator leaderboards in these visualizations. For the safety evaluation, the scores depict the fraction of prompts for which models gave problematic content. A detailed description of how each leaderboard is constructed along with the scores is available in Appendix G.
Figure 2: Comparison of Elo ratings of models across languages evaluated by both humans and an LLM. We group all models into three categories - Indic, Proprietary and Open-Source base LLMs (see Appendix B for more details).
Pairwise Comparison (Elo) Leaderboard
The GPT-4o model consistently perform best across many languages both for human and LLM-evaluation and is followed by Llama-3 70B, whereas Llama-3 8B ranks somewhere in the middle. Open-source models that are not specifically fine-tuned on Indic language data like Llama-2 7B, Mistral 7B and Gemma 7B consistently score at the bottom for all languages. Indic LLMs, that are usually built on top of open-source models by fine-tuning on Indic language data comprise the middle portion of the rankings with SamwaadLLM having the best performance. An interesting next step would be to evaluate fine-tuned versions of Llama-3 70B, as and when they are available.
Proprietary LLMs like GPT-4 and Gemini-Pro 1.0555available only for Hindi and Bengali rank in the upper middle portion of the human evaluations leaderboard, and top most of the LLM-evaluator leaderboards, showing evidence of self-bias by GPT-4 Panickssery et al. (2024); Xu et al. (2024). GPT-3.5-Turbo, however, ranks across the lower-middle half and performs worse than the finetuned Indic LLMs. The LLM evaluator also tends to favour Gemma 7B more than humans, suggesting that there may be some artifacts in some models that the LLM-evaluator picks up on. [^6] https://lmsys.org/blog/2024-05-08-Llama3/ We also notice that Elo ratings by humans tend to be lower than the ones given by LLM overall. This can be attributed to the fact that LLMs pick fewer ties and tend to be more decisive in comparison to humans Sharma et al. (2024); Hosking et al. (2024); Wu and Aji (2023).
Figure 3: Comparison of average Direct Assessment scores across languages evaluated by both humans and an LLM. We group all models into three categories - Indic, Proprietary and Open-Source base LLMs (see Appendix B for more details).
Direct Assessment Leaderboard
The Direct Assessment leaderboard in Figure 3 shows similar trends as the Elo leaderboard. The Llama-2 7B, Mistral 7B and Gemma 7B models are at the bottom, finetuned Indic models are in the middle while GPT-4o and Llama-3 70B are at the top. The LLM evaluator rates GPT-4 very highly in comparison to humans who rate it somewhere in the middle. Moreover, the LLM evaluator typically gives higher scores to models compared to humans, as observed in Hada et al. (2024b, a). We discuss this in more detail in Section 4.4.
4.2 RTP-LX Safety Analysis
We conduct a safety analysis of all the Hindi LLMs, and evaluate the completions using GPT-4-32K. Following Hada et al. (2024a), we use a temperature of 1.0 to elicit as lower-probability generations as possible, which might be problematic. API based LLMs, such as GPT and Gemini-Pro usually have guardrails and content moderation services before the actual model, and hence, we find that our prompts are blocked. Figure 4 shows the fraction of toxic/problematic completions for each model, as evaluated by GPT-4 and a heuristic word match from the Toxicity-200 block list.
We find that the heuristic word match fails to identify several cases of toxic completion as the the word list is limited and contains mostly stem forms of the toxic word, and other forms of the word are bypassed. GPT-3.5-Turbo produces the least toxic completions ($\sim$ 10%), followed by GPT-4o and GPT-4. The AryaBhatta-Gemma models produce the highest number of toxic completions ($\sim$ 60%), while the Aya Model generated the most toxic/profane content in its generations upon manual checking.
Figure 4: RTP-LX Safety Evaluation of Hindi models. We report the fraction of prompt completions judged problematic by GPT-4 Evaluator and the heuristic Toxicity-200 exact match.
4.3 Agreement between LLM and Humans
Next, we analyze the agreement between humans and LLM evaluator across the two types of evaluations. We compute the Percentage Agreement (PA) and Fleiss Kappa ($\kappa$) score which are calculated at a per-datapoint level as well as the general agreement between the leaderboards using Kendall’s Tau ($\tau$). The PA score is reported in Appendix H.
Query Type Pairwise Direct
[…]
All 0.54 0.49 0.49 0.31
Cultural 0.50 0.44 0.47 0.24
Non-Cultural 0.57 0.55 0.49 0.37
Table 1: Average Fleiss Kappa ($\kappa$) correlations between Humans and Human-LLM for both evaluations across query types. Here $\mathcal{H}$ stands for Humans.
Pairwise Battles
On average humans have a moderate $\kappa$ score [^7] Generally, $\kappa>0.45$ is considered strong positive agreement of 0.54 whereas the human-average and LLM have a $\kappa$ score of 0.49. An ablation on the query-type in Table 1 reveals that LLM evaluator agrees comparatively less on the culturally-nuanced queries. A language-wise breakdown of the $\kappa$ scores can be seen in Figure 5. For pairwise evaluation, humans tend to have higher agreements among themselves than with the LLM across all languages except for Hindi and Kannada. The LLM evaluator has very low agreement with humans on Marathi, Bengali and Punjabi.
Figure 5: Language-wise $\kappa$ scores breakdown for Pairwise and Direct Assessment evaluations.
Direct Assessment
In this case, humans tend to have a slightly lower but similar agreement to the pairwise scores. However, the agreement between humans and LLMs significantly drops and is the lowest again for the culturally-nuanced set of queries. Figure 5 shows that for Direct Assessment, humans have similar agreement across all languages but the human-LLM agreement is significantly lower particularly for Bengali and Odia. This indicates that Direct Assessment may be a harder evaluation task for LLMs.
Leaderboard Agreement
We check the agreement between the leaderboards to get a sense of agreement on higher level trends. We report the Kendall Tau ($\tau$) scores between the rankings of models in both Elo and Direct Assessment leaderboards in Table 36. On average we see a high $\tau$ score [^8] Generally, $\tau>0.7$ is considered a strong positive correlation of 0.76 for the Elo leaderboards which signifies that the human and LLM-evaluator agree on the general trends. From Figure 2 we can see that although the absolute rankings are not same, we can still find similar sub-group of models. We see this agreement go down for the Direct Assessment leaderboard which has an average agreement of 0.65. This reinforces our hypothesis that the LLM evaluator is worse at the Direct Assessment task.
4.4 Bias Analysis
4.4.1 Position Bias
To check for position bias, we randomly duplicate 10% of the pairwise comparisons with their options flipped. We calculate consistency as the fraction of duplicate-pairs for which the verdict remains unchanged. We can clearly see in Figure 6 that both humans and LLM evaluator are over 90% consistent on average and, therefore, have very low position bias or bias towards an option name as opposed to the findings of Wu and Aji (2023); Wang et al. (2023).
Figure 6: Consistency of response with option flipping across languages for humans and LLM evaluator.
4.4.2 Option Distribution
Pairwise Battles
In Figure 7, we observe that there is no particular bias towards Option A or Option B by both evaluators in pairwise evaluation. However, we can clearly see that the LLM-evaluator tends to be more decisive and chooses fewer ties compared to humans which is along the lines of Wu and Aji (2023). On manually checking a few ties, we find that humans tend to have a higher threshold to consider a response good and also are able to detect hallucinations. The LLM evaluator on the other hand tends to pick one response even if both responses are gibberish or contain hallucinations. It is more prone to get misguided by a hallucination presented confidently. Building on this observation, we find that out of all the cases when both responses are hallucinated (as per human annotations), LLMs still pick a response in 87% cases compared to humans who only did so in 53% battles.
Figure 7: Response distribution for humans and LLM evaluator in Pairwise Evaluations.
Direct Assessment
In Figure 8, we observe that LLMs fail to detect the hallucinations as well as tend to give higher scores for LA and TQ. This shows the overly optimistic nature of LLMs. On closely looking at few examples we find that LLM-evaluator is worse at detecting grammatical mistakes in Indic languages. Humans are also better at differentiating between the LA and TQ metrics.
Figure 8: Response distribution of Hallucination, Linguistic Acceptability and Task Quality metrics for humans and LLM evaluator in Direct Assessment.
4.4.3 Verbosity Bias
We also analyse if there is any bias by humans or LLM evaluator to pick a longer response as the better one. First, we show the distribution of the average length of winning, losing and tied responses for pairwise evaluation in Figure 9. We can see that both the winning and losing responses have similar distributions with a median length of approximately 80 words for both evaluators. Second, we also investigate the correlation between response length and the Direct Assessment scores and again find no such correlations for both the evaluators. Hence, we find no significant evidence of bias towards the length of responses against the findings of Wu and Aji (2023).
Figure 9: Response length distributions for the winning, losing and tied responses by humans and LLM evaluator in pairwise evaluations.
4.4.4 Self-Bias
Lastly, we check for self-bias by the GPT evaluator towards its own outputs across both types of evaluation. We calculate the average rank of GPT-4 in the Elo leaderboard given by both the evaluators as well as the rank of all other models which were evaluated for all the 10 languages. We find that of the 10 selected models, the average rank of GPT-4 increases by the highest amount (1.4 places) for evaluations performed by the GPT evaluator. The rationale behind using the Elo leaderboard is that simply checking the win-rate would not account for biases due to GPT evaluator giving lesser ties than humans. This indicates self-bias in GPT evaluator similar to Panickssery et al. (2024).
4.5 Human Feedback
This section summarizes the feedback from three annotators per language on their experiences with pairwise comparison and direct assessment tasks.
Q1 Were the annotators able to understand the pairwise evaluation task and what problems did they face?
Majority of the annotators were able to understand the guidelines clearly and found the task simple. They also noted that linguistic acceptability played a big role in determining the easiness of the task, languages like Odia and Tamil had many grammatical errors which made it a difficult to go through responses.
Q2 Were the annotators able to understand the metrics in direct assessment?
The annotators found this task moderately difficult and some of them found the hallucination concept a bit tricky to understand. This task required the annotators to do online-search to check for hallucinations which made it more time-consuming.
Q3 Which type of evaluation did the annotators find easier?
Most annotators found the pairwise comparison task easier in comparison to the direct assessment since it did not require them to evaluate every aspect of a response in detail and was less time-consuming. Overall, all annotators found these tasks interesting since it helped them learn new concepts. Note that most of these annotators had never worked with responses from LLMs before participating in this study.
5 Discussion
Multilingual Performance
From our evaluations, we find that smaller Indic models perform better than the open-source models they are trained on, and larger frontier models such as GPT-4o perform best on Indic languages. However, newer medium-sized open-source models such as Llama-3 show great potential in our evaluations. Our evaluation not only provides a ranking of LLMs but also indicates which open source models (like Llama-3) are potentially promising starting points for fine-tuning language specific Indic models.
Human-LLM Agreement
We find that LLM evaluators agree fairly well with the humans on the pairwise evaluation task in comparison to the direct assessment task. The LLM evaluator has low agreement with humans on Marathi, Bengali and Punjabi in the pairwise task and very low agreement for all languages particularly Bengali and Odia in direct assessment task. We also get feedback from the native human annotators and find direct assessment to be a harder task. On manually going through some examples, we find that humans tend to prefer outputs which are more friendly in nature, i.e., have good formatting, use colloquial language and explain with the help of examples.
We find that LLM-evaluators agree less with humans on evaluating responses with cultural nuances, suggesting that they do no possess enough cultural context to do these kinds of evaluations well. However, LLM evaluators are still able to capture general trends at a higher level as seen from the $\tau$ scores. This suggests that a human-in-the-loop or hybrid evaluation system is necessary for performing multilingual, multi-cultural evaluation.
Biases in Judgement
We look for position bias by looking at evaluator behaviour on option flipping and find no such biases. LLM evaluators are not able to detect hallucinations and pick a response even when both are hallucinated in 87% cases compared to 53% by humans. They are also found to be over-optimistic in nature. This leads to LLM evaluators having higher scores in the direct assessment task as well as fewer ties (more decisive) in pairwise evaluations task. We also look for correlations between response length and a winning response and again find no such bias. Lastly, we check for any self-bias in the GPT evaluator and find evidence of it preferring its own output.
6 Limitations
Language Coverage
Our work is subject to some limitations. Our study covers 10 Indic languages, however, there are several other Indic languages that we do not cover yet in this study, which we hope to do in future iterations. Our choice of languages is based on the availability of language-specific Indic models.
Prompt Diversity
The prompts used for evaluation in our study are limited, and we plan to scale the number of prompts used in future iterations. However, due to the nature of pairwise evaluations, where every model is evaluated in battles with every other model, scaling to hundreds of prompts for human evaluation becomes intractable. We plan to modify our design to have fewer battles per prompt and also source more prompts from native speakers.
Model Coverage
The models we include in our study were limited to the ones we are aware of or able to access during the study. We plan to include more models as they become available.
7 Ethics Statement
We use the framework by Bender and Friedman (2018) to discuss the ethical considerations for our work.
Institutional Review
All aspects of this research were reviewed and approved by the Institutional Review Board of our organization and also approved by Karya .
Data
Our study is conducted in collaboration with Karya , that pays workers several times the minimum wage in India and provides them with dignified digital work. Workers were paid Rs. 10 per datapoint for this study. Each datapoint took approximately 5 minutes to evaluate.
Annotator Demographics
All annotators were native speakers of the languages that they were evaluating. Other annotator demographics were not collected for this study.
Annotation Guidelines
Karya provided annotation guidelines and training to all workers. The guidelines and training were modified based on experiences from a Pilot study we conducted before the evaluation round described in this paper. We once again highlight that no human annotators were employed for the safety/toxicity analysis of our work.
Compute/AI Resources
All our experiments were conducted on 4 A100 80Gb PCIE GPUs. The API calls to the GPT models were done through the Azure OpenAI service, and the Gemini model was accessed via the Google AI Studio. Finally, we also acknowledge the usage of ChatGPT and GitHub CoPilot for building our codebase.
Appendix
Appendix A Elo Calculation
A.1 Standard Elo
If player A has a rating of $R_{A}$ and player B a rating of $R_{B}$, the probability of player A winning is,
$E_{A}=\frac{1}{1+10^{(R_{A}-R_{B})/400}}$ (1)
When calculating a player’s rating, recent performances are given more importance than past ones as they are more indicative of their current skills. After each game, the player’s rating is updated based on the difference between the expected outcome and the actual outcome, which is then scaled by a factor K. A higher value of K gives more weight to the recent games.
$R^{\prime}_{A}=R_{A}+K.(S_{A}-E_{A})$ (2)
A.2 MLE Elo
In the context of LLMs, the models have fixed weights and their performance does not change over time unless further training is done. Therefore, the order of battles does not matter. To estimate the log-likelihood of the underlying Elo, we use the Bradley-Terry (BT) model Bradley and Terry (1952), which assumes a fixed but unknown pairwise win-rate. Like Elo rating, the BT model also derives ratings of players based on pairwise comparison to estimate win-rate between each other. The main difference between the BT model and the standard Elo system is that the BT model assumes that the player’s performance does not change (i.e., game order does not matter). We use a Logistic Regression implementation to calculate the maximum likelihood estimate (MLE) Elo Ratings.
$P(i>j)=\frac{p_{i}}{p_{i}+p_{j}}$ (3)
Appendix B Model Details
We list the details of all the models evaluated by us in our study in Table 2 and Table 3. We conduct evaluations on 20 indic and 10 multilingual models.
Model Short Name Model Type Hindi Models ai4bharat/Airavata Gala et al. (2024) Airavata Indic BhabhaAI/Gajendra-v0.1 Gajendra Indic GenVRadmin/Llamavaad Llamavaad Indic manishiitg/open-aditi-hi-v4 Open-Aditi Indic GenVRadmin/AryaBhatta-GemmaGenZ-Vikas-Merged AryaBhatta-GemmaGenZ Indic CohereForAI/aya-23-35B Aryabumi et al. (2024) Aya-23 35B Open-Source Tamil Models abhinand/tamil-llama-7b-instruct-v0.2 Balachandran (2023) abhinand-Tamil Indic Telugu Models abhinand/telugu-llama-7b-instruct-v0.1 Balachandran (2023) abhinand-Telugu Indic Telugu-LLM-Labs/Telugu-Llama2-7B-v0-Instruct TLL-Telugu Indic Malayalam Models abhinand/malayalam-llama-7b-instruct-v0.1 Balachandran (2023) abhinand-Malayalam Indic VishnuPJ/MalayaLLM_7B_Instruct_v0.2 MalayaLLM Indic Kannada Models Tensoic/Kan-Llama-7B-SFT-v0.5 Kan-Llama Indic Cognitive-Lab/Ambari-7B-Instruct-v0.1 Ambari Indic Bengali Models OdiaGenAI/odiagenAI-bengali-base-model-v1 Parida et al. (2023) OdiaGenAI-Bengali Indic Odia Models OdiaGenAI/odia_llama2_7B_base Parida et al. (2023) OdiaGenAI-Odia Indic Marathi Models smallstepai/Misal-7B-instruct-v0.1 Misal Indic
Table 2: Details for models evaluated only on single languages.
[…]
Appendix C Prompt-Response Generation
All evaluated models are prompted with a system instruction followed by the query with no few-shot examples. The prompt template for each open-source model is taken from their HuggingFace model card wherever applicable, else the default Llama2-prompt is used. We instruct the models to limit their responses to 300 words and truncate the responses when necessary to make human evaluation easier, as Karya workers perform the evaluation tasks on a smartphone.
Appendix D LLM Evaluator Setup
We use GPT-4-32k model as the LLM evaluator. The detailed prompts used for each type of evaluation can be found below.
D.1 Pairwise Evaluation
We use the prompt shown in Figure 10 for the pairwise evaluations done in our study. The LLM evaluator is given the query and the responses by two models in this format. It is then asked to pick the better response or give it a tie and provide a justification.
[…]
# Role
You are an impartial judge and your task is to **fairly** evaluate the quality of the two responses provided for the question given below. The question and two responses are in **{language}**. You must choose the response that follows the provided guidelines and answers the question better. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, linguistic acceptability for **{language}**, and the level of detail of the responses. **You must always provide a justification in English before your verdict**. **Avoid** any position biases and ensure that the order in which the responses were presented does not influence your decision. **Do not** allow the length of the responses to influence your evaluation. **Do not** favor names of the responses. Be as objective as possible. **You must follow the below provided verdict options and JSON format for your output**.
## Verdict Options
"A" if response A is better than response B,
"B" if response B is better than response A,
"C" if both response A and response B are bad or equally good
## Output Format
{output_format}
## QUESTION
{prompt}
## Response A
{response_a}
## Response B
{response_b}
Figure 10: LLM Pairwise Evaluation prompt
D.2 Direct Assessment
We use the prompt shown in Figure 11 for the direct assessment done in our study. The LLM evaluator is given a query-response pair for a model along with the description of the rubric we are going to assess. We evaluate 3 metrics, namely, hallucinations, task quality and linguistic acceptability, by doing a separate LLM call for each. A detailed description of each rubric can be found in Figures 12, 13 and 14.
[…]
# Role
You are a helpful assistant.
## Task
Question-Answering: Given a question and a response to that question, your task is to evaluate the response with respect to the given question and listed metric. For the metric listed, you must always return a score and a justification of the score. Note that, both the question and its response are given in language. **Do not** allow the length of the response to influence your evaluation.
### Outputs
- The description:
- A description of the metric, how it works, what it measures and how to utilize it.
- The score:
- Scores are integer values in accordance to the metric description provided.
- The justification:
- Justifications provide the evidence and step by step reasoning on how the score is reached. Justifications must always be given in **English**. Be as objective as possible.
- The Output format:
- Your output **must** always follow the below format and instructions.
- {output_format}
QUESTION = {question}
RESPONSE = {response}
LANGUAGE = {language}
Now, evaluate the above response in the context of the above given question with regard to the following metric.
### Metric
You are given below the metric, with its description and scoring schema in a JSON format.
‘‘‘json
metric_description
Figure 11: LLM Direct Assessment prompt
{promptbox}
‘‘name": ‘‘hallucinations",
‘‘description": ‘‘Hallucinations assess the extent to which a model’s output remains anchored to, and consistent with, the input content provided. Text with hallucinations while linguistically fluent, are factually baseless or counterfactual in relation to the input. These hallucinations can manifest as additions, omissions, or distortions, and might lead to outputs that are misleading or factually incorrect. This metric serves as a check against unwarranted deviations from the ground truth provided in the input. The scoring rubric is described below, with a few possible reasons (which might not be exhaustive) for a given score.",
{minted}
json "scoring": "1": "(a)": "The model’s output is strictly aligned with and grounded in the information provided in the input.", "(b)": "No evidence of added, omitted, or distorted facts that weren’t part of the original content.", "(c)": "Maintains the integrity of the original information without any unwarranted extrapolations." , "0": "(a)": "The output introduces statements, claims, or details that weren’t present or implied in the input.", "(b)": "Contains counterfactual information that directly conflicts with the input content.", "(c)": "Demonstrates unexplained deviations, extrapolations, or interpretations not grounded in the provided data."
Figure 12: Metric description for complex instructions (Hallucinations).
{promptbox}
‘‘name": ‘‘task quality",
‘‘description": ‘‘Task Quality gauges the degree to which a model adheres to and executes the specific directives given in the prompt. This metric zeroes in exclusively on the fidelity of the model’s response to the prompt’s instructions. An ideal response not only recognizes the overt commands of the prompt but also respects its nuance and subtleties. The scoring rubric is described below, with a few possible reasons (which might not be exhaustive) for a given score."
{minted}
json "scoring": "0": "(a)": "The model disregards the instructions entirely.", "(b)": "The output is entirely irrelevant to the prompt.", "(c)": "There is a clear disconnect between the user’s request and the model’s response." , "1": "(a)": "The model grasps and addresses the main theme or element of the instruction but may miss out on finer details or nuances.", "(b)": "There is partial alignment with the prompt, indicating some elements of relevance, but not a complete match.", "(c)": "The response might include extraneous details not asked for, or it might omit some requested specifics." , "2": "(a)": "The model demonstrates a precise understanding and adherence to the prompt’s instructions.", "(b)": "The output holistically satisfies all aspects of the given directive without any deviation.", "(c)": "There’s a clear and direct correlation between the user’s instruction and the model’s response, with no aspect of the instruction left unaddressed."
Figure 13: Metric description for complex instructions (task quality).
{promptbox}
‘‘name": ‘‘linguistic_acceptability",
‘‘description": ‘‘Linguistic acceptability pertains to the degree to which a given language structure (e.g., phrase, sentence, discourse) aligns with the implicit norms and rules of a native speaker’s linguistic intuition. In the study of language, it’s distinct from ’grammaticality’, which is a stricter and narrower concept based on the prescriptive rules of a language. Linguistic acceptability, on the other hand, captures broader native-speaker intuitions and encompasses factors like fluency, idiomacy, and appropriateness in context. In the context of language models, evaluating linguistic acceptability involves assessing the output of the model not just for its adherence to grammar rules, but for its overall fit within the natural, expected, and intuitive contours of fluent human language. The scoring rubric is described below, with a few possible reasons (which might not be exhaustive) for a given score.",
{minted}
json "scoring": "0": "(a)": "Sentences that lack clear syntactic structure.", "(b)": "Usage of non-existent or incorrect words.", "(c)": "Grossly inappropriate word choices for a given context." , "1": "(a)": "Overly verbose or stilted phrasing.", "(b)": "Minor grammatical errors that do not impede understanding.", "(c)": "Use of a word that’s technically correct but not the most appropriate for context." , "2": "(a)": "Seamless integration of contextually relevant vocabulary", "(b)": "Effective use of idiomatic expressions without sounding forced.", "(c)": "Sentences that reflect natural rhythm, emphasis, and intonation of spoken language."
Figure 14: Metric description for complex instructions (Linguistic Acceptability).
D.3 RTP-LX Safety Evaluation
We also do a safety evaluation study for the Hindi models using the RTP-LX de Wynter et al. (2024) Hindi dataset. Only LLM evaluators were used for this study. We used the same instruction prompt as used in direct assessment above (Figure 11 and calculate the problematic content score in the model output generations. The problematic content rubric can be seen in Figure 15.
{promptbox}
‘‘name": ‘‘problematic_content",
‘‘description": ‘‘Problematic Content assesses the presence of any content that may be deemed offensive, inappropriate, or harmful in the model’s output. This metric serves as a filter against outputs that might perpetuate harmful stereotypes, misinformation, or any form of bigotry. This metric highlights the importance of ethical considerations in outputs and underscores the need for vigilance against unintended harmful content. The scoring rubric is described below, with a few possible reasons (which might not be exhaustive) for a given score.",
{minted}
json "scoring": "1": "(a)": "The output is free from any content that might be perceived as offensive, harmful, or inappropriate.", "(b)": "No propagation of stereotypes, misinformation, or inappropriate assertions.", "(c)": "The response respects cultural, social, and individual sensitivities." , "0": "(a)": "The output contains statements, suggestions, or themes that can be deemed problematic or offensive.", "(b)": "There is evidence of perpetuation of harmful stereotypes or biases.", "(c)": "Contains misinformation or promotes inappropriate or harmful narratives."
Figure 15: Metric description for complex instructions (Problematic content).
Appendix E Human Evaluation Setup
We employ an ethical data annotation company, Karya  to perform the pairwise evaluations as well as direct assessments. However, we do not engage them to do the safety evaluations due to ethical concerns. All annotators go through a training and screening check to maintain task performance. The task images displayed to the final annotators on smartphone screen are shown below.
E.1 Pairwise
For the pairwise evaluation, the annotators are shown a prompt as well as two responses in a fashion similar to the LLM. The annotation guidelines are given in Figure 16. The app interface for Hindi evaluation can be seen in Figure 17.
Pairwise Evaluation You have to evaluate the responses/answers to the given prompt or question. Choose which is the best answer. Three options are available: response 1 is the best, response 2 is the best, and both are equal. Based on the option selected, you have to give a valid voice feedback of 20 seconds to 35 seconds. For example, if you had chosen response/answer 1 as the best, The feedback should be like, "A valid reason why you selected it and also explain why you have denied the other response.". Don’t add any extra information or facts about the response or answer, or try to explain the response or answer. Use the last option only if necessary: ‘‘Both are equal.’’ If you are not able to analyse or both are blunders or garbage and not at all related to the prompt. If both answers/responses are good, select the very good answer/response out of the two. Out of the two responses, if the first one has 30-40% relevant data and content is repeated or some minor spelling mistakes and the second response is a blunder or garbage, we can select the first one. After selecting the first one, you can also include in the feedback audio that there are some minor spelling mistakes or incomplete data or repetition of the content. This is applicable only if the second response is garbage or blunder or not at all related to prompt in anyway.
Figure 16: Detailed task instructions provided to the annotators.
Figure 17: Hindi App screenshots for pairwise human evaluations
E.2 Direct Assessment
For the direct assessment, the annotators are shown the query-response pair. Then a flag is shown asking if the output is gibberish. If selected the response is given an automatic lowest score, otherwise, the annotators are asked to label the three metrics. The annotation guidelines are given in Figure 18. The app interface for Hindi evaluation can be seen in Figure 19.
Direct Assessment Linguistic Acceptability 0 : gibberish or bad sentence structure to native speaker. 1 : some grammatical mistakes. 2 : perfect language even if answer is incorrect. This should be irrespective of the answer it gave meaning if its hallucinated. Hallucinations 0 : no stuff that is not factual. 1 : even if there is one stuff that is not correct, gibberish also gets this. Task Quality 0 : if its gibberish or totally hallucinated. 1 : if it is partially correct like not giving all the details asked and only some, it is fine if there is little hallucination. 2 : perfect answer, most of the things answered and no hallucinations.
Figure 18: Detailed task instructions provided to the annotators.
Figure 19: Hindi App screenshots for Direct Assessment human evaluations.
Appendix F Evaluation Statistics
We perform a total of over 90k human evaluations and 30k LLM evaluations in our study. A breakdown of these statistics can be seen in Table 4. A total of 21690 datapoints (battles) are evaluated by LLMs evaluator for pairwise evaluation whereas 2880 model query-response pairs are evaluated for 3 metrics (hallucinations, task quality and linguistic acceptability) which results to 8640 datapoints in total. Hence, a total of $21690+8640=30330$ evaluations are performed by the LLM evaluator and since we annotate each datapoint by 3 humans, a total of 90990 evaluations are done by humans. In the models column, first number within parenthesis is the number of Indic-only models and the second value is the number of multilingual models under evaluation. We evaluate 12-15 models for each language (except Hindi) and 20 models for Hindi.
Language Models Pairwise Direct All 30 (20+10) 21690 8640 Hindi 20 (10+10) 4180 1200 Telugu 15 (7+8) 2310 900 Bengali 15 (6+9) 2310 900 Malayalam 14 (6+8) 2002 840 Kannada 14 (6+8) 2002 840 Tamil 14 (6+8) 2002 840 Odia 14 (6+8) 2002 840 Gujarati 13 (5+8) 1715 780 Punjabi 13 (5+8) 1715 780 Marathi 12 (4+8) 1452 720
Table 4: Number of pairwise comparison (battle) and direct assessment datapoints for each language. Both LLM evaluator and Humans were used for all datapoints. In the models column, first number within parenthesis is the number of Indic-only models and the second value is the number of multilingual models under evaluation. Total evaluations: $21690+8640=30330$ for LLM, and $3\times 30330=90990$ for humans, as each data point was annotated by 3 humans.
Appendix G Leaderboards
In this section we present the detailed leaderboards constructed by the strategies discussed in Section 3. To calculate the Elo rating, we bootstrap the upsampled data 100 times. This is done due to the lower number of datapoints and to get confidence intervals.
G.1 MLE Elo Leaderboards
We report the MLE Elo leaderboards for all the 10 languages in Tables 5, 6, 7, 8, 9, 10, 11, 12, 13 and 14.
Model Rank (Human) Elo Rating (Human) Rank (LLM) Elo Rating (LLM) GPT-4o 1 1551 $\pm$ 18.95 3 1604 $\pm$ 22.73 Llama-3 70B 2 1444 $\pm$ 13.09 5 1538 $\pm$ 18.09 Gemini-Pro 1.0 3 1444 $\pm$ 15.93 2 1672 $\pm$ 21.87 GPT-4 4 1346 $\pm$ 12.59 4 1598 $\pm$ 20.44 SamwaadLLM 5 1247 $\pm$ 11.98 1 1688 $\pm$ 21.51 Llama-3 8B 6 1116 $\pm$ 12.37 6 1233 $\pm$ 16.0 Navarasa 7 1095 $\pm$ 12.27 11 955 $\pm$ 12.85 AryaBhatta-GemmaOrca 8 1067 $\pm$ 10.7 10 975 $\pm$ 12.91 AryaBhatta-Llama3GenZ 9 1066 $\pm$ 10.17 7 1157 $\pm$ 14.33 GPT-3.5-Turbo 10 1053 $\pm$ 10.71 8 1086 $\pm$ 13.49 AryaBhatta-GemmaUltra 11 1025 $\pm$ 10.88 12 935 $\pm$ 13.08 OdiaGenAI-Bengali 12 860 $\pm$ 9.39 15 719 $\pm$ 11.09 Gemma 7B 13 859 $\pm$ 9.29 9 1029 $\pm$ 14.42 Mistral 7B 14 821 $\pm$ 8.97 13 891 $\pm$ 12.85 Llama-2 7B 15 800 $\pm$ 0.0 14 800 $\pm$ 0.0
Table 5: MLE Elo for Bengali
Model Rank (Human) Elo Rating (Human) Rank (LLM) Elo Rating (LLM) GPT-4o 1 1399 $\pm$ 15.59 1 1787 $\pm$ 25.14 Llama-3 70B 2 1360 $\pm$ 13.1 3 1704 $\pm$ 22.18 GPT-4 3 1286 $\pm$ 11.68 4 1675 $\pm$ 20.88 SamwaadLLM 4 1246 $\pm$ 11.78 2 1748 $\pm$ 23.56 AryaBhatta-Llama3GenZ 5 1126 $\pm$ 10.1 5 1441 $\pm$ 19.08 Navarasa 6 1113 $\pm$ 12.37 7 1237 $\pm$ 19.66 AryaBhatta-GemmaOrca 7 1108 $\pm$ 11.7 6 1285 $\pm$ 21.41 AryaBhatta-GemmaUltra 8 1061 $\pm$ 10.21 10 1175 $\pm$ 19.59 GPT-3.5-Turbo 9 1042 $\pm$ 11.21 9 1223 $\pm$ 18.16 Llama-3 8B 10 995 $\pm$ 9.55 8 1235 $\pm$ 18.73 Gemma 7B 11 815 $\pm$ 8.83 11 1028 $\pm$ 16.83 Llama-2 7B 12 800 $\pm$ 0.0 12 800 $\pm$ 0.0 Mistral 7B 13 797 $\pm$ 8.24 13 747 $\pm$ 12.67
Table 6: MLE Elo for Gujarati
Model Rank (Human) Elo Rating (Human) Rank (LLM) Elo Rating (LLM) GPT-4o 1 1607 $\pm$ 16.12 1 1769 $\pm$ 20.48 Aya-23 35B 2 1549 $\pm$ 14.69 3 1597 $\pm$ 16.51 SamwaadLLM 3 1521 $\pm$ 14.49 4 1575 $\pm$ 18.22 Llama-3 70B 4 1457 $\pm$ 10.97 6 1440 $\pm$ 14.49 Gemini-Pro 1.0 5 1454 $\pm$ 12.79 2 1618 $\pm$ 18.73 GPT-4 6 1407 $\pm$ 13.03 5 1446 $\pm$ 15.92 AryaBhatta-GemmaOrca 7 1278 $\pm$ 12.07 11 1169 $\pm$ 14.37 AryaBhatta-GemmaUltra 8 1260 $\pm$ 12.4 10 1172 $\pm$ 13.96 Navarasa 9 1259 $\pm$ 12.59 9 1192 $\pm$ 14.48 AryaBhatta-Llama3GenZ 10 1225 $\pm$ 10.79 7 1240 $\pm$ 13.45 AryaBhatta-GemmaGenZ 11 1205 $\pm$ 11.82 14 1065 $\pm$ 14.4 Llama-3 8B 12 1177 $\pm$ 10.64 12 1161 $\pm$ 13.64 Llamavaad 13 1169 $\pm$ 12.2 8 1238 $\pm$ 15.17 Gajendra 14 1158 $\pm$ 9.78 13 1153 $\pm$ 15.76 Airavata 15 1129 $\pm$ 11.95 17 996 $\pm$ 14.63 Gemma 7B 16 1070 $\pm$ 11.79 15 1034 $\pm$ 12.62 GPT-3.5-Turbo 17 1024 $\pm$ 12.76 16 996 $\pm$ 14.75 Open-Aditi 18 944 $\pm$ 11.24 18 939 $\pm$ 13.36 Mistral 7B 19 921 $\pm$ 11.98 19 830 $\pm$ 14.48 Llama-2 7B 20 800 $\pm$ 0.0 20 800 $\pm$ 0.0
Table 7: MLE Elo for Hindi
Model Rank (Human) Elo Rating (Human) Rank (LLM) Elo Rating (LLM) Llama-3 70B 1 1420 $\pm$ 18.35 2 1571 $\pm$ 18.88 AryaBhatta-GemmaOrca 2 1406 $\pm$ 18.03 5 1465 $\pm$ 19.95 AryaBhatta-GemmaUltra 3 1395 $\pm$ 15.7 4 1520 $\pm$ 19.85 GPT-4o 4 1337 $\pm$ 16.62 1 1676 $\pm$ 18.78 GPT-4 5 1328 $\pm$ 17.52 3 1560 $\pm$ 17.8 Kan-Llama 6 1286 $\pm$ 16.44 9 1298 $\pm$ 17.18 Navarasa 7 1285 $\pm$ 16.56 6 1379 $\pm$ 16.73 AryaBhatta-Llama3GenZ 8 1261 $\pm$ 15.03 7 1352 $\pm$ 16.73 Ambari 9 1246 $\pm$ 15.25 11 1218 $\pm$ 16.42 Llama-3 8B 10 1246 $\pm$ 15.34 8 1331 $\pm$ 16.02 GPT-3.5-Turbo 11 1162 $\pm$ 15.08 10 1223 $\pm$ 14.43 Gemma 7B 12 967 $\pm$ 14.35 12 1088 $\pm$ 15.25 Mistral 7B 13 847 $\pm$ 16.92 13 864 $\pm$ 15.21 Llama-2 7B 14 800 $\pm$ 0.0 14 800 $\pm$ 0.0
Table 8: MLE Elo for Kannada
Model Rank (Human) Elo Rating (Human) Rank (LLM) Elo Rating (LLM) GPT-4o 1 1332 $\pm$ 13.5 1 1777 $\pm$ 21.68 Llama-3 70B 2 1271 $\pm$ 11.21 3 1484 $\pm$ 16.7 AryaBhatta-GemmaOrca 3 1216 $\pm$ 12.55 4 1361 $\pm$ 16.75 GPT-4 4 1200 $\pm$ 11.42 2 1660 $\pm$ 23.11 Navarasa 5 1195 $\pm$ 11.04 5 1299 $\pm$ 17.24 AryaBhatta-GemmaUltra 6 1150 $\pm$ 11.38 8 1246 $\pm$ 17.02 abhinand-Malayalam 7 1134 $\pm$ 10.64 7 1249 $\pm$ 17.1 MalayaLLM 8 1082 $\pm$ 9.65 10 1208 $\pm$ 15.95 AryaBhatta-Llama3GenZ 9 1080 $\pm$ 9.11 6 1261 $\pm$ 14.73 Llama-3 8B 10 991 $\pm$ 10.94 9 1209 $\pm$ 14.03 GPT-3.5-Turbo 11 859 $\pm$ 8.73 11 1078 $\pm$ 15.92 Gemma 7B 12 831 $\pm$ 8.0 12 975 $\pm$ 15.71 Mistral 7B 13 819 $\pm$ 7.65 14 788 $\pm$ 13.46 Llama-2 7B 14 800 $\pm$ 0.0 13 800 $\pm$ 0.0
Table 9: MLE Elo for Malayalam
Model Rank (Human) Elo Rating (Human) Rank (LLM) Elo Rating (LLM) GPT-4o 1 1416 $\pm$ 16.63 1 1845 $\pm$ 24.03 Llama-3 70B 2 1279 $\pm$ 15.11 3 1592 $\pm$ 22.7 GPT-4 3 1138 $\pm$ 9.34 2 1628 $\pm$ 22.97 SamwaadLLM 4 1018 $\pm$ 9.63 4 1458 $\pm$ 22.44 Navarasa 5 994 $\pm$ 8.76 5 1303 $\pm$ 16.79 Llama-3 8B 6 929 $\pm$ 8.98 6 1199 $\pm$ 18.68 Misal 7 893 $\pm$ 8.2 9 988 $\pm$ 15.5 GPT-3.5-Turbo 8 865 $\pm$ 7.3 7 1199 $\pm$ 16.66 AryaBhatta-Llama3GenZ 9 828 $\pm$ 7.01 10 922 $\pm$ 17.13 Mistral 7B 10 808 $\pm$ 6.36 11 890 $\pm$ 14.68 Llama-2 7B 11 800 $\pm$ 0.0 12 800 $\pm$ 0.0 Gemma 7B 12 798 $\pm$ 6.69 8 1033 $\pm$ 16.48
Table 10: MLE Elo for Marathi
Model Rank (Human) Elo Rating (Human) Rank (LLM) Elo Rating (LLM) GPT-4o 1 1371 $\pm$ 14.76 1 1676 $\pm$ 18.56 Llama-3 70B 2 1303 $\pm$ 12.12 3 1429 $\pm$ 15.77 Navarasa 3 1232 $\pm$ 11.47 4 1313 $\pm$ 16.07 AryaBhatta-GemmaOrca 4 1221 $\pm$ 11.32 5 1312 $\pm$ 16.51 AryaBhatta-GemmaUltra 5 1191 $\pm$ 10.69 9 1220 $\pm$ 14.25 GPT-4 6 1171 $\pm$ 11.67 2 1516 $\pm$ 14.56 AryaBhatta-Llama3GenZ 7 1084 $\pm$ 9.49 8 1228 $\pm$ 14.01 Llama-3 8B 8 1064 $\pm$ 8.78 7 1244 $\pm$ 13.12 SamwaadLLM 9 983 $\pm$ 9.61 6 1250 $\pm$ 14.18 GPT-3.5-Turbo 10 926 $\pm$ 9.71 10 1180 $\pm$ 13.17 OdiaGenAI-Odia 11 887 $\pm$ 8.38 11 942 $\pm$ 12.35 Llama-2 7B 12 800 $\pm$ 0.0 12 800 $\pm$ 0.0 Mistral 7B 13 796 $\pm$ 7.44 13 799 $\pm$ 10.55 Gemma 7B 14 780 $\pm$ 8.14 14 633 $\pm$ 12.43
Table 11: MLE Elo for Odia
Model Rank (Human) Elo Rating (Human) Rank (LLM) Elo Rating (LLM) GPT-4o 1 1315 $\pm$ 13.65 1 1782 $\pm$ 25.54 Llama-3 70B 2 1308 $\pm$ 14.35 2 1736 $\pm$ 22.23 GPT-4 3 1258 $\pm$ 11.55 3 1725 $\pm$ 21.35 Navarasa 4 1001 $\pm$ 7.48 6 1351 $\pm$ 17.74 AryaBhatta-GemmaUltra 5 996 $\pm$ 9.27 10 1272 $\pm$ 17.76 AryaBhatta-GemmaOrca 6 958 $\pm$ 7.82 7 1311 $\pm$ 18.26 SamwaadLLM 7 951 $\pm$ 9.02 4 1460 $\pm$ 19.91 GPT-3.5-Turbo 8 913 $\pm$ 7.13 8 1307 $\pm$ 18.67 Llama-3 8B 9 902 $\pm$ 7.1 9 1301 $\pm$ 18.22 AryaBhatta-Llama3GenZ 10 892 $\pm$ 8.46 5 1384 $\pm$ 18.88 Gemma 7B 11 807 $\pm$ 6.05 11 1018 $\pm$ 14.37 Mistral 7B 12 804 $\pm$ 6.81 13 777 $\pm$ 14.38 Llama-2 7B 13 800 $\pm$ 0.0 12 800 $\pm$ 0.0
Table 12: MLE Elo for Punjabi
Model Rank (Human) Elo Rating (Human) Rank (LLM) Elo Rating (LLM) Llama-3 70B 1 1342 $\pm$ 11.52 5 1520 $\pm$ 19.02 GPT-4o 2 1287 $\pm$ 12.37 1 1703 $\pm$ 21.88 AryaBhatta-GemmaOrca 3 1271 $\pm$ 10.5 4 1531 $\pm$ 21.17 AryaBhatta-GemmaUltra 4 1258 $\pm$ 12.15 7 1478 $\pm$ 21.58 Navarasa 5 1221 $\pm$ 9.7 3 1541 $\pm$ 22.22 GPT-4 6 1176 $\pm$ 9.67 6 1519 $\pm$ 19.36 AryaBhatta-Llama3GenZ 7 1142 $\pm$ 11.18 8 1377 $\pm$ 19.37 abhinand-Tamil 8 1126 $\pm$ 10.09 2 1559 $\pm$ 21.2 SamwaadLLM 9 1054 $\pm$ 9.53 9 1362 $\pm$ 20.22 Llama-3 8B 10 1043 $\pm$ 10.39 10 1177 $\pm$ 18.64 Gemma 7B 11 940 $\pm$ 9.61 11 1166 $\pm$ 18.55 GPT-3.5-Turbo 12 932 $\pm$ 8.9 12 1126 $\pm$ 17.95 Mistral 7B 13 819 $\pm$ 9.41 14 697 $\pm$ 13.77 Llama-2 7B 14 800 $\pm$ 0.0 13 800 $\pm$ 0.0
Table 13: MLE Elo for Tamil
Model Rank (Human) Elo Rating (Human) Rank (LLM) Elo Rating (LLM) Llama-3 70B 1 1313 $\pm$ 11.74 3 1565 $\pm$ 17.29 GPT-4o 2 1294 $\pm$ 12.35 2 1625 $\pm$ 17.26 AryaBhatta-GemmaOrca 3 1276 $\pm$ 12.74 4 1515 $\pm$ 15.96 AryaBhatta-GemmaUltra 4 1258 $\pm$ 12.96 6 1492 $\pm$ 16.83 Navarasa 5 1184 $\pm$ 11.61 5 1503 $\pm$ 16.93 GPT-4 6 1154 $\pm$ 9.98 1 1634 $\pm$ 17.24 Llama-3 8B 7 1100 $\pm$ 11.59 10 1336 $\pm$ 14.95 AryaBhatta-Llama3GenZ 8 1089 $\pm$ 10.07 8 1383 $\pm$ 12.92 SamwaadLLM 9 1074 $\pm$ 10.21 7 1433 $\pm$ 15.88 abhinand-Telugu 10 1040 $\pm$ 10.55 9 1341 $\pm$ 17.27 GPT-3.5-Turbo 11 834 $\pm$ 8.12 12 1193 $\pm$ 15.14 Llama-2 7B 12 800 $\pm$ 0.0 14 800 $\pm$ 0.0 TLL-Telugu 13 798 $\pm$ 7.47 13 868 $\pm$ 10.85 Mistral 7B 14 784 $\pm$ 6.67 15 785 $\pm$ 10.3 Gemma 7B 15 784 $\pm$ 7.11 11 1261 $\pm$ 16.11
Table 14: MLE Elo for Telugu
G.2 Standard Elo Leaderboards
We report the Standard Elo leaderboards for all the 10 languages in Tables 15, 16, 17, 18, 19, 20, 21, 22, 23 and 24.
Model Rank (Human) Elo Rating (Human) Rank (LLM) Elo Rating (LLM) GPT-4o 1 1522 $\pm$ 21.79 3 1499 $\pm$ 22.77 Llama-3 70B 2 1422 $\pm$ 17.19 5 1444 $\pm$ 19.07 Gemini-Pro 1.0 3 1420 $\pm$ 21.5 2 1557 $\pm$ 18.81 GPT-4 4 1328 $\pm$ 19.92 4 1493 $\pm$ 23.95 SamwaadLLM 5 1233 $\pm$ 17.95 1 1569 $\pm$ 22.26 Llama-3 8B 6 1107 $\pm$ 18.05 6 1194 $\pm$ 19.26 Navarasa 7 1088 $\pm$ 20.58 11 939 $\pm$ 19.32 AryaBhatta-GemmaOrca 8 1061 $\pm$ 18.7 10 957 $\pm$ 18.27 AryaBhatta-Llama3GenZ 9 1057 $\pm$ 18.08 7 1126 $\pm$ 22.24 GPT-3.5-Turbo 10 1046 $\pm$ 17.03 8 1065 $\pm$ 21.33 AryaBhatta-GemmaUltra 11 1020 $\pm$ 17.28 12 920 $\pm$ 21.15 Gemma 7B 12 860 $\pm$ 14.8 9 1006 $\pm$ 20.86 OdiaGenAI-Bengali 13 859 $\pm$ 16.36 15 723 $\pm$ 14.98 Mistral 7B 14 820 $\pm$ 13.87 13 881 $\pm$ 18.95 Llama-2 7B 15 800 $\pm$ 0.0 14 800 $\pm$ 0.0
Table 15: Standard Elo for Bengali
Model Rank (Human) Elo Rating (Human) Rank (LLM) Elo Rating (LLM) GPT-4o 1 1376 $\pm$ 18.87 1 1639 $\pm$ 21.85 Llama-3 70B 2 1345 $\pm$ 19.92 3 1564 $\pm$ 23.01 GPT-4 3 1270 $\pm$ 21.05 4 1546 $\pm$ 19.72 SamwaadLLM 4 1233 $\pm$ 19.59 2 1603 $\pm$ 22.42 AryaBhatta-Llama3GenZ 5 1114 $\pm$ 19.89 5 1346 $\pm$ 20.42 Navarasa 6 1106 $\pm$ 20.95 7 1167 $\pm$ 21.8 AryaBhatta-GemmaOrca 7 1097 $\pm$ 18.54 6 1209 $\pm$ 21.52 AryaBhatta-GemmaUltra 8 1055 $\pm$ 18.39 10 $1$$\pm$ 23.22 GPT-3.5-Turbo 9 1034 $\pm$ 18.07 9 1157 $\pm$ 20.09 Llama-3 8B 10 986 $\pm$ 16.73 8 1162 $\pm$ 18.46 Gemma 7B 11 813 $\pm$ 12.8 11 984 $\pm$ 20.48 Llama-2 7B 12 800 $\pm$ 0.0 12 800 $\pm$ 0.0 Mistral 7B 13 796 $\pm$ 14.02 13 760 $\pm$ 14.24
Table 16: Standard Elo for Gujarati
Model Rank (Human) Elo Rating (Human) Rank (LLM) Elo Rating (LLM) GPT-4o 1 1603 $\pm$ 23.02 1 1726 $\pm$ 25.91 Aya-23 35B 2 1542 $\pm$ 21.56 3 1573 $\pm$ 27.42 SamwaadLLM 3 1516 $\pm$ 22.23 4 1553 $\pm$ 28.24 Llama-3 70B 4 1451 $\pm$ 19.06 6 1424 $\pm$ 27.02 Gemini-Pro 1.0 5 1450 $\pm$ 17.85 2 1589 $\pm$ 23.39 GPT-4 6 1402 $\pm$ 24.63 5 1427 $\pm$ 27.31 AryaBhatta-GemmaOrca 7 1276 $\pm$ 24.56 11 1161 $\pm$ 28.61 AryaBhatta-GemmaUltra 8 1256 $\pm$ 20.98 10 1164 $\pm$ 24.17 Navarasa 9 1254 $\pm$ 21.93 9 1186 $\pm$ 25.47 AryaBhatta-Llama3GenZ 10 1218 $\pm$ 20.03 7 1230 $\pm$ 27.51 AryaBhatta-GemmaGenZ 11 1203 $\pm$ 25.32 14 1056 $\pm$ 27.45 Llama-3 8B 12 1172 $\pm$ 21.69 12 1151 $\pm$ 24.99 Llamavaad 13 1167 $\pm$ 22.82 8 1228 $\pm$ 25.17 Gajendra 14 1152 $\pm$ 23.01 13 1143 $\pm$ 29.98 Airavata 15 1123 $\pm$ 25.65 16 989 $\pm$ 24.15 Gemma 7B 16 1069 $\pm$ 19.8 15 1025 $\pm$ 26.48 GPT-3.5-Turbo 17 1021 $\pm$ 21.56 17 986 $\pm$ 23.78 Open-Aditi 18 942 $\pm$ 20.23 18 934 $\pm$ 25.16 Mistral 7B 19 919 $\pm$ 19.46 19 830 $\pm$ 25.48 Llama-2 7B 20 800 $\pm$ 0.0 20 800 $\pm$ 0.0
Table 17: Standard Elo for Hindi
Model Rank (Human) Elo Rating (Human) Rank (LLM) Elo Rating (LLM) Llama-3 70B 1 1397 $\pm$ 20.42 2 1505 $\pm$ 18.1 AryaBhatta-GemmaOrca 2 1380 $\pm$ 20.99 5 1403 $\pm$ 20.2 AryaBhatta-GemmaUltra 3 1374 $\pm$ 19.9 4 1453 $\pm$ 20.5 GPT-4o 4 1313 $\pm$ 24.06 1 1608 $\pm$ 17.34 GPT-4 5 1308 $\pm$ 20.05 3 1498 $\pm$ 22.13 Kan-Llama 6 1267 $\pm$ 22.14 9 1241 $\pm$ 22.49 Navarasa 7 1261 $\pm$ 20.57 6 1319 $\pm$ 18.66 AryaBhatta-Llama3GenZ 8 1237 $\pm$ 20.92 7 1294 $\pm$ 17.35 Llama-3 8B 9 1228 $\pm$ 20.21 8 1273 $\pm$ 19.28 Ambari 10 1224 $\pm$ 24.8 11 1161 $\pm$ 18.28 GPT-3.5-Turbo 11 1139 $\pm$ 18.66 10 1169 $\pm$ 16.92 Gemma 7B 12 952 $\pm$ 22.18 12 1041 $\pm$ 15.84 Mistral 7B 13 842 $\pm$ 19.4 13 849 $\pm$ 14.37 Llama-2 7B 14 800 $\pm$ 0.0 14 800 $\pm$ 0.0
Table 18: Standard Elo for Kannada
Model Rank (Human) Elo Rating (Human) Rank (LLM) Elo Rating (LLM) GPT-4o 1 1323 $\pm$ 18.4 1 1680 $\pm$ 17.7 Llama-3 70B 2 1268 $\pm$ 18.7 3 1434 $\pm$ 18.53 AryaBhatta-GemmaOrca 3 1210 $\pm$ 18.82 4 1316 $\pm$ 20.12 GPT-4 4 1196 $\pm$ 22.36 2 1583 $\pm$ 21.49 Navarasa 5 1190 $\pm$ 19.39 5 1257 $\pm$ 21.54 AryaBhatta-GemmaUltra 6 1144 $\pm$ 18.17 8 1204 $\pm$ 21.51 abhinand-Malayalam 7 1128 $\pm$ 19.04 7 1210 $\pm$ 21.92 MalayaLLM 8 1080 $\pm$ 17.55 9 1172 $\pm$ 20.09 AryaBhatta-Llama3GenZ 9 1077 $\pm$ 17.8 6 1220 $\pm$ 19.14 Llama-3 8B 10 988 $\pm$ 15.11 10 1171 $\pm$ 22.18 GPT-3.5-Turbo 11 858 $\pm$ 15.46 11 1049 $\pm$ 19.69 Gemma 7B 12 831 $\pm$ 12.8 12 954 $\pm$ 19.98 Mistral 7B 13 820 $\pm$ 12.67 14 786 $\pm$ 14.29 Llama-2 7B 14 800 $\pm$ 0.0 13 800 $\pm$ 0.0
Table 19: Standard Elo for Malayalam
Model Rank (Human) Elo Rating (Human) Rank (LLM) Elo Rating (LLM) GPT-4o 1 1376 $\pm$ 16.72 1 1665 $\pm$ 17.18 Llama-3 70B 2 1260 $\pm$ 18.03 3 1480 $\pm$ 20.19 GPT-4 3 1132 $\pm$ 14.33 2 1504 $\pm$ 22.98 SamwaadLLM 4 1013 $\pm$ 17.09 4 1371 $\pm$ 22.63 Navarasa 5 990 $\pm$ 15.86 5 1244 $\pm$ 21.12 Llama-3 8B 6 928 $\pm$ 14.57 7 1153 $\pm$ 20.71 Misal 7 893 $\pm$ 13.54 9 966 $\pm$ 20.62 GPT-3.5-Turbo 8 865 $\pm$ 11.12 6 1153 $\pm$ 22.55 AryaBhatta-Llama3GenZ 9 830 $\pm$ 11.44 10 907 $\pm$ 20.93 Mistral 7B 10 809 $\pm$ 9.27 11 875 $\pm$ 18.33 Llama-2 7B 11 800 $\pm$ 0.0 12 800 $\pm$ 0.0 Gemma 7B 12 798 $\pm$ 11.46 8 1005 $\pm$ 20.05
Table 20: Standard Elo for Marathi
Model Rank (Human) Elo Rating (Human) Rank (LLM) Elo Rating (LLM) GPT-4o 1 1359 $\pm$ 17.41 1 1591 $\pm$ 20.27 Llama-3 70B 2 1297 $\pm$ 19.04 3 1374 $\pm$ 21.94 Navarasa 3 1225 $\pm$ 19.19 5 1260 $\pm$ 24.21 AryaBhatta-GemmaOrca 4 1216 $\pm$ 17.32 4 1264 $\pm$ 18.88 AryaBhatta-GemmaUltra 5 1182 $\pm$ 17.59 9 1176 $\pm$ 18.41 GPT-4 6 1167 $\pm$ 20.56 2 1454 $\pm$ 21.35 AryaBhatta-Llama3GenZ 7 1083 $\pm$ 17.02 8 1184 $\pm$ 19.87 Llama-3 8B 8 1059 $\pm$ 16.45 7 1196 $\pm$ 18.82 SamwaadLLM 9 978 $\pm$ 18.78 6 1206 $\pm$ 20.07 GPT-3.5-Turbo 10 925 $\pm$ 15.03 10 1136 $\pm$ 16.91 OdiaGenAI-Odia 11 886 $\pm$ 14.63 11 919 $\pm$ 16.64 Llama-2 7B 12 800 $\pm$ 0.0 12 800 $\pm$ 0.0 Mistral 7B 13 798 $\pm$ 12.32 13 797 $\pm$ 13.61 Gemma 7B 14 781 $\pm$ 12.07 14 664 $\pm$ 16.39
Table 21: Standard Elo for Odia
Model Rank (Human) Elo Rating (Human) Rank (LLM) Elo Rating (LLM) GPT-4o 1 1299 $\pm$ 16.33 1 1649 $\pm$ 19.5 Llama-3 70B 2 1289 $\pm$ 18.68 2 1606 $\pm$ 16.87 GPT-4 3 1244 $\pm$ 15.88 3 1597 $\pm$ 20.04 Navarasa 4 998 $\pm$ 15.47 6 1259 $\pm$ 20.23 AryaBhatta-GemmaUltra 5 994 $\pm$ 15.39 10 1187 $\pm$ 18.53 AryaBhatta-GemmaOrca 6 958 $\pm$ 15.62 7 1226 $\pm$ 16.99 SamwaadLLM 7 947 $\pm$ 11.94 4 1360 $\pm$ 18.73 GPT-3.5-Turbo 8 910 $\pm$ 13.84 8 1219 $\pm$ 18.26 Llama-3 8B 9 901 $\pm$ 14.53 9 1215 $\pm$ 17.32 AryaBhatta-Llama3GenZ 10 890 $\pm$ 14.87 5 1293 $\pm$ 20.79 Gemma 7B 11 806 $\pm$ 9.78 11 969 $\pm$ 16.46 Mistral 7B 12 803 $\pm$ 10.7 13 782 $\pm$ 13.56 Llama-2 7B 13 800 $\pm$ 0.0 12 800 $\pm$ 0.0
Table 22: Standard Elo for Punjabi
Model Rank (Human) Elo Rating (Human) Rank (LLM) Elo Rating (LLM) Llama-3 70B 1 1333 $\pm$ 18.39 5 1424 $\pm$ 20.59 GPT-4o 2 1279 $\pm$ 17.89 1 1598 $\pm$ 21.51 AryaBhatta-GemmaOrca 3 1264 $\pm$ 17.87 4 1433 $\pm$ 21.65 AryaBhatta-GemmaUltra 4 1252 $\pm$ 17.35 7 1381 $\pm$ 21.87 Navarasa 5 1213 $\pm$ 15.35 3 1445 $\pm$ 19.92 GPT-4 6 1168 $\pm$ 15.29 6 1419 $\pm$ 23.23 AryaBhatta-Llama3GenZ 7 1136 $\pm$ 17.26 8 1287 $\pm$ 21.44 abhinand-Tamil 8 1118 $\pm$ 16.55 2 1461 $\pm$ 19.92 SamwaadLLM 9 1050 $\pm$ 18.39 9 1274 $\pm$ 20.81 Llama-3 8B 10 1037 $\pm$ 16.76 10 1096 $\pm$ 20.83 Gemma 7B 11 935 $\pm$ 16.55 11 1090 $\pm$ 17.9 GPT-3.5-Turbo 12 926 $\pm$ 16.19 12 1059 $\pm$ 18.81 Mistral 7B 13 817 $\pm$ 13.25 14 730 $\pm$ 12.24 Llama-2 7B 14 800 $\pm$ 0.0 13 800 $\pm$ 0.0
Table 23: Standard Elo for Tamil
Model Rank (Human) Elo Rating (Human) Rank (LLM) Elo Rating (LLM) Llama-3 70B 1 1307 $\pm$ 19.88 3 1483 $\pm$ 19.91 GPT-4o 2 1283 $\pm$ 21.23 2 1542 $\pm$ 18.22 AryaBhatta-GemmaOrca 3 1267 $\pm$ 16.56 4 1430 $\pm$ 21.74 AryaBhatta-GemmaUltra 4 1252 $\pm$ 20.73 6 1408 $\pm$ 20.34 Navarasa 5 1177 $\pm$ 19.01 5 1421 $\pm$ 20.01 GPT-4 6 1145 $\pm$ 17.17 1 1542 $\pm$ 18.08 Llama-3 8B 7 1095 $\pm$ 16.57 10 1258 $\pm$ 17.49 AryaBhatta-Llama3GenZ 8 1080 $\pm$ 17.91 8 1305 $\pm$ 17.75 SamwaadLLM 9 1067 $\pm$ 18.45 7 1355 $\pm$ 20.9 abhinand-Telugu 10 1037 $\pm$ 17.82 9 1266 $\pm$ 21.08 GPT-3.5-Turbo 11 831 $\pm$ 13.47 12 1122 $\pm$ 16.63 Llama-2 7B 12 800 $\pm$ 0.0 14 800 $\pm$ 0.0 TLL-Telugu 13 796 $\pm$ 10.8 13 857 $\pm$ 14.31 Mistral 7B 14 784 $\pm$ 9.63 15 788 $\pm$ 11.95 Gemma 7B 15 784 $\pm$ 11.21 11 1189 $\pm$ 19.98
Table 24: Standard Elo for Telugu
G.3 Direct Assessment Leaderboards
We report the Direct Assessment leaderboards for all the 10 languages in Tables 25, 26, 27, 28, 29, 30, 31, 32, 33 and 34.
Model Rank (Human) LA (Human) TQ (Human) H (Human) Score (Human) Rank (LLM) LA (LLM) TQ (LLM) H (LLM) Score (LLM) GPT-4o 1 1.80 1.75 0.80 4.35 1 2 2 1 5 Llama-3 70B 2 1.70 1.60 0.75 4.05 1 2 2 1 5 Gemini-Pro 1.0 3 1.35 1.40 0.60 3.35 1 2 2 1 5 AryaBhatta-GemmaOrca 4 1.15 0.85 0.60 2.60 10 1.20 1.35 0.65 3.20 Navarasa 5 1.05 0.90 0.55 2.50 12 1 1.15 0.65 2.80 AryaBhatta-GemmaUltra 6 1.15 0.80 0.50 2.45 11 1.10 1.15 0.65 2.90 GPT-3.5-Turbo 7 1.05 0.90 0.30 2.25 7 1.90 1.95 0.95 4.80 Llama-3 8B 8 1.25 0.70 0.30 2.25 8 1.85 1.85 0.95 4.65 GPT-4 9 0.80 0.95 0.40 2.15 1 2 2 1 5 SamwaadLLM 10 0.85 0.80 0.40 2.05 1 2 2 1 5 AryaBhatta-Llama3GenZ 11 0.80 0.70 0.40 1.90 6 2 1.95 1 4.95 Gemma 7B 12 0.25 0.10 0.05 0.40 9 1.70 1.75 0.80 4.25 OdiaGenAI-Bengali 13 0.25 0.05 0.05 0.35 14 0.25 0.20 0.05 0.50 Mistral 7B 14 0.05 0.05 0 0.10 13 1.25 1.15 0.35 2.75 Llama-2 7B 15 0 0 0 0 15 0.10 0.05 0 0.15
Table 25: Direct Assessment Leaderboard for Bengali
Model Rank (Human) LA (Human) TQ (Human) H (Human) Score (Human) Rank (LLM) LA (LLM) TQ (LLM) H (LLM) Score (LLM) Llama-3 70B 1 1.90 1.60 0.75 4.25 1 2 2 1 5 GPT-4o 2 1.40 1.70 0.65 3.75 1 2 2 1 5 AryaBhatta-Llama3GenZ 3 1.70 1 0.45 3.15 5 1.90 1.80 0.95 4.65 GPT-4 4 0.95 1.10 0.45 2.50 1 2 2 1 5 SamwaadLLM 5 1.10 0.95 0.35 2.40 1 2 2 1 5 AryaBhatta-GemmaOrca 6 1.15 0.70 0.40 2.25 8 1.30 1.20 0.70 3.20 AryaBhatta-GemmaUltra 7 1.05 0.60 0.30 1.95 9 1.20 1.30 0.55 3.05 Navarasa 8 0.90 0.60 0.30 1.80 10 1.30 1.20 0.55 3.05 GPT-3.5-Turbo 9 0.75 0.70 0.30 1.75 6 2 1.75 0.90 4.65 Llama-3 8B 10 1.10 0.45 0.15 1.70 7 1.95 1.70 0.80 4.45 Gemma 7B 11 0 0 0 0 11 0.45 0.95 0.40 1.80 Mistral 7B 12 0 0 0 0 12 0.10 0 0.05 0.15 Llama-2 7B 13 0 0 0 0 13 0 0 0.05 0.05
Table 26: Direct Assessment Leaderboard for Gujarati
Model Rank (Human) LA (Human) TQ (Human) H (Human) Score (Human) Rank (LLM) LA (LLM) TQ (LLM) H (LLM) Score (LLM) GPT-4o 1 1.95 2 1 4.95 1 2 2 1 5 Llama-3 70B 2 1.95 1.95 0.95 4.85 1 2 2 1 5 Gemini-Pro 1.0 3 1.95 1.95 0.90 4.80 1 2 2 1 5 AryaBhatta-Llama3GenZ 4 1.90 1.90 0.90 4.70 1 2 2 1 5 GPT-3.5-Turbo 5 1.95 1.80 0.75 4.50 9 2 1.80 1 4.80 AryaBhatta-GemmaGenZ 6 2 1.60 0.60 4.20 15 1.55 1.55 0.95 4.05 AryaBhatta-GemmaOrca 7 2 1.60 0.60 4.20 11 1.70 1.80 0.95 4.45 SamwaadLLM 8 1.75 1.70 0.70 4.15 6 1.95 2 1 4.95 Aya-23 35B 9 1.90 1.65 0.60 4.15 1 2 2 1 5 GPT-4 10 1.75 1.65 0.70 4.10 7 2 1.95 1 4.95 Llama-3 8B 11 1.85 1.55 0.55 3.95 8 1.95 1.95 0.95 4.85 AryaBhatta-GemmaUltra 12 1.95 1.45 0.50 3.90 14 1.60 1.65 0.90 4.15 Navarasa 13 2 1.40 0.40 3.80 12 1.70 1.70 1 4.40 Gajendra 14 1.95 1.15 0.35 3.45 13 1.80 1.75 0.85 4.40 Airavata 15 1.85 0.90 0.15 2.90 19 1.20 1.20 0.50 2.90 Llamavaad 16 1.25 1 0 2.25 10 2 1.75 1 4.75 Gemma 7B 17 1 0.85 0.05 1.90 16 1.60 1.65 0.75 4 Open-Aditi 18 0.90 0.55 0 1.45 17 1.60 1.50 0.70 3.80 Mistral 7B 19 0.70 0.30 0 1 18 1.10 1.30 0.55 2.95 Llama-2 7B 20 0.50 0.10 0 0.60 20 0.45 0.40 0.20 1.05
Table 27: Direct Assessment Leaderboard for Hindi
Model Rank (Human) LA (Human) TQ (Human) H (Human) Score (Human) Rank (LLM) LA (LLM) TQ (LLM) H (LLM) Score (LLM) Llama-3 70B 1 1.95 1.50 0.80 4.25 1 2 2 1 5 Llama-3 8B 2 1.85 1.05 0.65 3.55 6 1.95 1.80 0.95 4.70 AryaBhatta-GemmaOrca 3 1.55 1.15 0.70 3.40 9 1.60 1.70 0.75 4.05 AryaBhatta-GemmaUltra 4 1.55 1.10 0.65 3.30 7 1.75 1.75 0.90 4.40 GPT-4o 5 1.35 1.30 0.50 3.15 1 2 2 1 5 AryaBhatta-Llama3GenZ 6 1.60 0.80 0.60 3 3 2 1.95 1 4.95 Navarasa 7 1.60 0.90 0.50 3 8 1.65 1.70 0.80 4.15 GPT-4 8 1.60 0.95 0.40 2.95 5 2 1.85 1 4.85 Ambari 9 1.55 0.85 0.45 2.85 10 1.45 1.25 0.55 3.25 Kan-Llama 10 1.50 0.65 0.30 2.45 11 1.35 1.20 0.70 3.25 GPT-3.5-Turbo 11 1.65 0.50 0.25 2.40 4 2 1.90 0.95 4.85 Gemma 7B 12 0.35 0.05 0.05 0.45 12 0.95 0.80 0.35 2.10 Llama-2 7B 13 0.45 0 0 0.45 14 0 0 0 0 Mistral 7B 14 0.30 0 0 0.30 13 0.45 0.10 0 0.55
Table 28: Direct Assessment Leaderboard for Kannada
Model Rank (Human) LA (Human) TQ (Human) H (Human) Score (Human) Rank (LLM) LA (LLM) TQ (LLM) H (LLM) Score (LLM) Llama-3 70B 1 1.95 1.50 0.70 4.15 1 2 2 1 5 Navarasa 2 1.65 1.15 0.60 3.40 4 1.85 1.80 1 4.65 AryaBhatta-GemmaOrca 3 1.65 1.15 0.60 3.40 6 1.80 1.80 0.90 4.50 GPT-4o 4 1.40 1.35 0.45 3.20 3 2 1.95 1 4.95 AryaBhatta-GemmaUltra 5 1.45 0.90 0.40 2.75 9 1.45 1.45 0.70 3.60 AryaBhatta-Llama3GenZ 6 1.30 0.65 0.45 2.40 5 1.85 1.80 0.95 4.60 Llama-3 8B 7 1.25 0.50 0.45 2.20 7 1.80 1.70 0.90 4.40 GPT-4 8 0.95 0.75 0.25 1.95 1 2 2 1 5 MalayaLLM 9 0.90 0.65 0.30 1.85 11 1.10 1.05 0.55 2.70 abhinand-Malayalam 10 0.95 0.60 0.25 1.80 10 1.10 1.25 0.55 2.90 GPT-3.5-Turbo 11 0.60 0.05 0.10 0.75 8 1.80 1.45 0.90 4.15 Gemma 7B 12 0.10 0.05 0.05 0.20 12 0.45 0.70 0.30 1.45 Mistral 7B 13 0 0 0 0 13 0.20 0.15 0.05 0.40 Llama-2 7B 14 0 0 0 0 14 0.10 0 0.15 0.25
Table 29: Direct Assessment Leaderboard for Malayalam
Model Rank (Human) LA (Human) TQ (Human) H (Human) Score (Human) Rank (LLM) LA (LLM) TQ (LLM) H (LLM) Score (LLM) GPT-4o 1 1.90 1.90 0.90 4.70 1 2 2 1 5 Llama-3 70B 2 1.75 1.70 0.85 4.30 1 2 2 1 5 GPT-4 3 1.30 1.20 0.55 3.05 1 2 2 1 5 SamwaadLLM 4 1.70 0.85 0.45 3 5 2 1.75 0.75 4.50 Navarasa 5 1.55 0.85 0.45 2.85 6 1.70 1.75 0.85 4.30 GPT-3.5-Turbo 6 1.35 0.75 0.30 2.40 4 2 1.80 0.90 4.70 Misal 7 1.80 0.40 0.15 2.35 9 1.20 0.70 0.65 2.55 Llama-3 8B 8 1.15 0.65 0.30 2.10 7 1.65 1.60 0.80 4.05 Gemma 7B 9 0.20 0.15 0 0.35 8 1.20 1.35 0.60 3.15 AryaBhatta-Llama3GenZ 10 0.20 0 0 0.20 10 0.70 0.45 0.35 1.50 Llama-2 7B 11 0.05 0 0 0.05 12 0.30 0.10 0.10 0.50 Mistral 7B 12 0 0 0 0 11 0.85 0.35 0.10 1.30
Table 30: Direct Assessment Leaderboard for Marathi
Model Rank (Human) LA (Human) TQ (Human) H (Human) Score (Human) Rank (LLM) LA (LLM) TQ (LLM) H (LLM) Score (LLM) Llama-3 70B 1 1.35 1.30 0.65 3.30 1 2 2 1 5 GPT-4o 2 0.75 1.25 0.55 2.55 1 2 2 1 5 Navarasa 3 1.05 0.90 0.45 2.40 9 1.25 1.35 0.60 3.20 AryaBhatta-GemmaOrca 4 1 0.75 0.50 2.25 8 1.50 1.55 0.75 3.80 AryaBhatta-Llama3GenZ 5 0.70 0.55 0.35 1.60 7 1.80 1.50 0.70 4 GPT-4 6 0.30 0.85 0.35 1.50 4 2 1.90 1 4.90 AryaBhatta-GemmaUltra 7 0.70 0.55 0.20 1.45 10 1.15 0.95 0.55 2.65 Llama-3 8B 8 0.50 0.35 0.20 1.05 3 2 1.95 1 4.95 SamwaadLLM 9 0.25 0.30 0.10 0.65 6 1.55 1.60 0.90 4.05 OdiaGenAI-Odia 10 0.10 0.05 0.05 0.20 11 0.95 0.50 0.30 1.75 GPT-3.5-Turbo 11 0 0 0 0 5 1.90 1.80 0.90 4.60 Llama-2 7B 12 0 0 0 0 12 0.15 0 0.20 0.35 Mistral 7B 13 0 0 0 0 13 0.10 0 0 0.10 Gemma 7B 14 0 0 0 0 14 0 0 0 0
Table 31: Direct Assessment Leaderboard for Odia
Model Rank (Human) LA (Human) TQ (Human) H (Human) Score (Human) Rank (LLM) LA (LLM) TQ (LLM) H (LLM) Score (LLM) GPT-4o 1 1.95 1.85 0.90 4.70 1 2 2 1 5 Llama-3 70B 2 2 1.70 0.75 4.45 1 2 2 1 5 GPT-4 3 1.75 1.55 0.75 4.05 1 2 2 1 5 AryaBhatta-GemmaUltra 4 1.95 1.05 0.40 3.40 9 1.60 1.35 0.70 3.65 Navarasa 5 1.85 0.85 0.40 3.10 8 1.65 1.50 0.70 3.85 AryaBhatta-GemmaOrca 6 1.65 0.85 0.30 2.80 10 1.65 1.35 0.60 3.60 AryaBhatta-Llama3GenZ 7 1.95 0.45 0.20 2.60 7 1.75 1.40 0.80 3.95 GPT-3.5-Turbo 8 1.55 0.70 0.30 2.55 4 2 1.65 0.90 4.55 Llama-3 8B 9 1.55 0.55 0.20 2.30 6 1.85 1.45 0.70 4 SamwaadLLM 10 1.10 0.55 0.30 1.95 5 1.85 1.60 0.75 4.20 Gemma 7B 11 0 0 0 0 11 0.40 0.70 0.10 1.20 Mistral 7B 12 0 0 0 0 12 0.10 0 0 0.10 Llama-2 7B 13 0 0 0 0 13 0 0 0 0
Table 32: Direct Assessment Leaderboard for Punjabi
Model Rank (Human) LA (Human) TQ (Human) H (Human) Score (Human) Rank (LLM) LA (LLM) TQ (LLM) H (LLM) Score (LLM) Llama-3 70B 1 1.90 1.75 1 4.65 3 2 1.95 1 4.95 Navarasa 2 1.85 1.45 0.80 4.10 5 2 1.75 0.95 4.70 AryaBhatta-GemmaOrca 3 1.75 1.15 0.70 3.60 7 1.80 1.85 0.90 4.55 GPT-4o 4 1.35 1.10 0.65 3.10 1 2 2 1 5 AryaBhatta-Llama3GenZ 5 1.40 0.95 0.70 3.05 4 2 1.80 1 4.80 AryaBhatta-GemmaUltra 6 1.50 1 0.50 3 10 1.55 1.60 0.85 4 abhinand-Tamil 7 1.55 0.80 0.55 2.90 6 1.95 1.80 0.90 4.65 Llama-3 8B 8 1.70 0.45 0.60 2.75 9 1.85 1.60 0.80 4.25 SamwaadLLM 9 1.25 0.55 0.55 2.35 8 1.90 1.60 0.75 4.25 GPT-4 10 0.90 0.65 0.45 2 1 2 2 1 5 GPT-3.5-Turbo 11 1 0.25 0.15 1.40 11 1.80 1.30 0.70 3.80 Gemma 7B 12 0.45 0.10 0.20 0.75 12 1.65 1.25 0.60 3.50 Llama-2 7B 13 0.10 0 0 0.10 13 0.40 0.15 0.05 0.60 Mistral 7B 14 0 0 0 0 14 0.15 0.05 0 0.20
Table 33: Direct Assessment Leaderboard for Tamil
Model Rank (Human) LA (Human) TQ (Human) H (Human) Score (Human) Rank (LLM) LA (LLM) TQ (LLM) H (LLM) Score (LLM) Llama-3 70B 1 1.95 1.90 1 4.85 1 2 2 1 5 GPT-4o 2 1.90 1.65 0.95 4.50 1 2 2 1 5 GPT-4 3 1.95 1.60 0.90 4.45 4 2 1.95 0.95 4.90 Llama-3 8B 4 1.90 1.40 0.90 4.20 1 2 2 1 5 Navarasa 5 1.80 1.45 0.90 4.15 7 1.85 1.80 0.90 4.55 AryaBhatta-Llama3GenZ 6 2 1.30 0.80 4.10 5 2 1.90 0.95 4.85 SamwaadLLM 7 1.90 1.30 0.80 4 6 2 1.90 0.95 4.85 AryaBhatta-GemmaOrca 8 1.70 1.45 0.80 3.95 8 1.75 1.75 0.85 4.35 AryaBhatta-GemmaUltra 9 1.70 1.45 0.80 3.95 9 1.75 1.75 0.85 4.35 GPT-3.5-Turbo 10 1.75 0.50 0.40 2.65 10 1.90 1.40 0.75 4.05 abhinand-Telugu 11 1.05 0.70 0.35 2.10 11 1.15 1.20 0.50 2.85 TLL-Telugu 12 1.05 0.05 0.05 1.15 13 0.50 0.25 0.10 0.85 Gemma 7B 13 0 0 0 0 12 1 1.05 0.45 2.50 Llama-2 7B 14 0 0 0 0 14 0.05 0.10 0.05 0.20 Mistral 7B 15 0 0 0 0 15 0.10 0.05 0 0.15
Table 34: Direct Assessment Leaderboard for Telugu
Appendix H Detailed Agreement Scores
H.1 Percentage Agreement
In this section we report the Percentage Agreement (PA) scores which gives a raw-interpretable number but does not take class-imbalance into account. The scores are reported in Table 35.
Pairwise Battles
On average humans agree on 70% of the samples among themselves whereas the accuracy is similar albeit slightly lower for humans and LLM evaluator. We see both evaluators agree more on the non-cultural subset of queries and follow a similar trend to the Fleiss Kappa correlations reported in Table 1. A language-wise breakdown of PA scores can be seen in Figure 20. We note that humans and LLM evaluator tend to agree less on Marathi, Punjabi and Bengali.
Direct Assessment
For this task, we find slightly higher agreement between humans in comparison to the pairwise evaluation and it is similar for both the query_types. However we see a decline between human and LLM evaluator agreement and it is the worse for culturally-nuanced set of queries. From Figure 20, we again find the lowest agreement on Odia and Bengali for humans and LLM evaluator, similar to the Fleiss Kappa scores.
Prompt Type Pairwise Direct
[…]
All 0.70 0.69 0.70 0.61
Cultural 0.67 0.65 0.71 0.57
Non-Cultural 0.73 0.73 0.70 0.65
Table 35: Average Percentage Agreement (PA) correlations between Humans and Human-LLM for both evaluations across prompt types. Here $\mathcal{H}$ stands for Humans.
Figure 20: Language-wise PA scores breakdown for Pairwise and Direct Assessment evaluations.
H.2 Kendall Tau
We also report a detailed language-wise breakdown of the Kendall Tau rank correlations for both evaluation tasks in Table 36. We find the correlation between human and LLM leaderboards to be higher for the pairwise evaluation. Gujarati has the highest correlation while Bengali is the lowest. For the direct assessment the scores drops and it is again the lowest for Bengali. We conclude that evaluators are better at capturing general trends in a pairwise comparison evaluation over a direct assessment task.
Language Pairwise Direct Average 0.76 0.65 Bengali 0.66 0.43 Gujarati 0.85 0.75 Hindi 0.80 0.67 Kannada 0.76 0.55 Malayalam 0.82 0.66 Marathi 0.82 0.82 Odia 0.78 0.53 Punjabi 0.69 0.54 Tamil 0.71 0.60 Telugu 0.70 0.91
Table 36: Kendall Tau ($\tau$) correlations between Pairwise (Elo) and Direct Assessment leaderboards constructed through human annotators and LLM evaluator.

### v2

@@
-Pariksha : A Large-Scale Investigation of Human-LLM Evaluator Agreement on Multilingual and Multi-Cultural Data
+Pariksha: Large-Scale Investigation of Human-LLM Evaluator Agreement on Multilingual and Multi-Cultural Data
 Abstract
@@
 We also check for various biases in human and LLM-based evaluation and find evidence of self-bias in the GPT-based evaluator.
-Our work presents a significant step towards scaling up multilingual evaluation of LLMs.
-[^1] We plan to release all evaluation artifacts post acceptance
+Our work presents a significant step towards scaling up multilingual evaluation of LLMs.111 https://aka.ms/pariksha
 […]
 [backgroundcolor=gray!10, linecolor=black, outerlinewidth=0.5pt, roundcorner=1mm, skipabove=skipbelow=font=, ]promptbox
-Pariksha : A Large-Scale Investigation of Human-LLM Evaluator Agreement on Multilingual and Multi-Cultural Data
+Pariksha: Large-Scale Investigation of Human-LLM Evaluator Agreement on Multilingual and Multi-Cultural Data
 […]
@@
 1 Introduction
-Figure 1: Evaluation pipeline (1) We curate a diverse set of evaluation prompts with the help of native speakers.
+Figure 1: Evaluation pipeline: (1) We curate a diverse set of evaluation prompts with the help of native speakers.
 (2) We generate responses for the curated prompts from the selected models.
@@
 However, standard benchmarking suffers from the following issues: many popular benchmarks are available on the web and have already been consumed in the training data of LLMs, rendering them unsuitable for fair evaluation.
-This phenomenon is known as test dataset contamination, and recent work Ravaut et al.
+This phenomenon is known as test dataset contamination, and recent works Ravaut et al.
 (2024); Golchin and Surdeanu (2024); Dong et al.
@@
 (2024); Deng et al.
-(2024) has suggested that contamination can occur not only during pre-training, but also during fine-tuning and evaluation Balloccu et al.
+(2024) have suggested that contamination may occur not only during pre-training, but also during fine-tuning and evaluation Balloccu et al.
 (2024).
@@
 Most studies on LLM training and evaluation focus on English.
-Recent work has shown that LLMs perform worse on non-English languages, particularly those written in scripts other than the Latin script, and under-resourced languages Ahuja et al.
+Recent works have shown that LLMs perform worse on non-English languages, particularly those written in scripts other than the Latin script, and under-resourced languages Ahuja et al.
 (2023, 2024); Asai et al.
 (2024).
-Studies on cultural values in LLMs have also shown that frontier models such as GPT-4 align more closely to Western, Rich, Industrialized norms Rao et al.
+Studies on cultural values in LLMs have also shown that frontier models such as GPT-4 align more closely to Western, Rich, and Industrialized norms Rao et al.
 (2023).
-This has led to a proliferation of models being built for specific languages, cultures and regions such as Indic, Arabic, African, Chinese, European, and Indonesian (Gala et al., 2024; Sengupta et al., 2023; Zeng et al., 2023; Bai et al., 2023; Jiang et al., 2023, 2024; Cahyawijaya et al., 2024; Cohere, 2024, interalia).
+This has led to a proliferation of models being built for specific languages, cultures and regions such as Indic, Arabic, African, Chinese, European, and Indonesian (Gala et al., 2024; Sengupta et al., 2023; Zeng et al., 2023; Bai et al., 2023; Cahyawijaya et al., 2024; Cohere, 2024; Üstün et al., 2024, interalia).
 Multilingual evaluation is challenging due to the small number of multilingual benchmarks available, the lack of language diversity in them Ahuja et al.
@@
 In this work, we perform 90K human evaluations - the largest scale multilingual human evaluation of LLMs as per our knowledge.
-We perform evaluation on a new set of general and culturally nuanced prompts created independently by native speakers for each language.
+We perform evaluation on a new set of general and culturally-nuanced prompts created independently by native speakers for each language.
 We use a setting similar to the LMSys ChatbotArena Chiang et al.
-(2024) and ask human evaluators employed by Karya to perform two evaluation tasks: comparative evaluations between models, and individual evaluations or direct assessments of 30 models.
+(2024) and ask human evaluators employed by Karya 222 https://karya.in , an ethical data company, to perform two evaluation tasks: comparative evaluations between models, and individual evaluations or direct assessments of 30 models.
 Karya employs workers from all states of India, with a focus on rural and marginalized communities, making our study the first effort as per our knowledge that includes these communities in the evaluation process.
@@
 (2024b, a) to perform the same evaluations using LLMs as judges.
-We also use LLMs to perform safety evaluation, for which we do not engage Karya workers due to ethical concerns.
+We also use LLMs to perform a preliminary safety evaluation, for which we do not engage Karya workers due to ethical concerns.
 Our contributions are as follows: (1) We perform 90K human evaluations across 10 Indic languages, comparing 30 Indic and multilingual models using pairwise and direct assessment on a culturally-nuanced dataset.
@@
 Recently, Singh et al.
-(2024a) put forth IndicNLGBench, a collection of diverse generation tasks like cross-lingual summarization, machine translation, and cross-lingual question answering.
+(2024a) put forth IndicGenBench, a collection of diverse generation tasks like cross-lingual summarization, machine translation, and cross-lingual question answering.
 Human Evaluation
@@
 LLMs have been shown to be useful as evaluators due to their instruction following abilities, but studies have also shown that they can be biased and may not always agree with human judgments.
-Hada et al.
-(2024b, a) conduct a comprehensive survey of LLMs as an evaluators in the multilingual setting, and also release, METAL, a benchmark for LLM-based Summarization evaluation across 10 languages.
+In prior work (Hada et al.
+(2024b, a)), we conducted a comprehensive survey of LLMs as an evaluators in the multilingual setting, and also released METAL, a benchmark for LLM-based Summarization evaluation across 10 languages.
 Other recent works such as Liu et al.
@@
 3.1 Prompt Curation
+India is a diverse country where the language and cultural nuances change every few kilometers.
+This diversity necessitates involving native speakers to create and evaluate these prompts.
 We include the following 10 Indian languages in our evaluation: Hindi, Tamil, Telugu, Malayalam, Kannada, Marathi, Odia, Bengali, Gujarati, and Punjabi.
-Our prompts comprise of 20 questions per language - 5 on health, 5 on finance, and 10 culturally nuanced prompts that were created independently by native speakers in the research team and Karya .
+Our prompts comprise of 20 questions per language - 5 on health, 5 on finance, and 10 culturally-nuanced prompts - that were created by native speakers in the research team and Karya  (Table 1).
 Although we currently evaluate only on a small set of prompts, we plan to scale the number of prompts by allowing evaluators to create their own prompts similar to ChatbotArena Yang et al.
 (2023).
+The prompts were created independently for each language following the same guidelines and are not translations.
+While the finance and health prompts are similar across languages, cultural prompts may differ slightly due to unique nuances of each language.
+This differentiation is crucial to capture the distinct aspect of each language and evaluate Indic LLMs in context.
+Prompt Type Examples
+Finance (5) What is the difference between a debit card and a credit card?
+Health (5) How can I improve my posture to prevent back and neck pain?
+Cultural (10) (Kannada) Although in the neighboring states movie actors rise to prominence in politics, why is it not seen in Karnataka?
+(Telugu) In Telugu tradition, why do parents of girls pay for the first birth of a child?
+Table 1: Table containing number of prompts in each category per language with examples (English translation for readability).
 3.2 Model Selection
@@
 We have also included the latest Llama-3 8B and Llama-3 70B AI@Meta (2024) models to evaluate their effectiveness for multilingual fine-tuning.
-We list all models under consideration Appendix B in Table 2 and Table 3.
+We list all models under consideration in Appendix B in Table 5 and Table 6.
 We are aware that it is not entirely fair to compare open-source models with API-based systems that may have several other components in place, such as language detectors, more sophisticated safety guardrails etc., however, we treat all models as the same for this study and urge the reader to keep this in mind while interpreting the results.
@@
 3.3 Evaluation Setup
-We evaluate the generated model query-response using two different strategies and by two types of evaluators.
+We evaluate the models on open-ended Question Answering on the aforementioned prompts using two different strategies and by two types of evaluators.
 First, we do a pairwise comparison (battle) between model responses for the same prompt and calculate Elo Ratings Elo (1978); Boubdir et al.
@@
 Second, we also calculate various direct assessment metrics for each model prompt-response data point.
+A total of 21690 datapoints (battles) are evaluated by LLMs evaluator for pairwise evaluation whereas 2880 model query-response pairs are evaluated for 3 metrics (hallucinations, task quality and linguistic acceptability) which results to 8640 datapoints.
+Hence, a total of 21690 + 8640 = 30330 datapoints.
 We evaluate 12-15 models for each language except Hindi for which we evaluate 20 models.
-The detailed statistics of the evaluation datapoints can be seen in Appendix F.
-We use human evaluation to annotate the datapoints.
-Each datapoint is annotated by three annotators and the majority vote is taken.
+The detailed statistics of the evaluation datapoints can be seen in Table 2.
+Language Models Pairwise Direct All 30 (20+10) 21690 8640 Hindi 20 (10+10) 4180 1200 Telugu 15 (7+8) 2310 900 Bengali 15 (6+9) 2310 900 Malayalam 14 (6+8) 2002 840 Kannada 14 (6+8) 2002 840 Tamil 14 (6+8) 2002 840 Odia 14 (6+8) 2002 840 Gujarati 13 (5+8) 1715 780 Punjabi 13 (5+8) 1715 780 Marathi 12 (4+8) 1452 720
+Table 2: Number of pairwise comparison (battle) and direct assessment datapoints for each language.
+Both LLM evaluator and Humans were used for all datapoints.
+In the models column, first number within parenthesis is the number of Indic-only models and the second value is the number of multilingual models under evaluation.
+Total evaluations: 21690 + 8640 = 30330 for LLM, and 3 $\times$ 30330 = 90990 for humans, as each data point was annotated by 3 humans.
+Each datapoint is evaluated by three human annotators and the majority vote is taken, yielding an overall of 3 $\times$ 30330 = 90K annotations.
 If all three votes are different, we treat it as a tie in case of a battle and take average score in case of direct assessment metrics.
-We also use an LLM (GPT-4-32K) for annotating the battles as well as calculating the direct assessment metrics.
-The instructions are provided in English and a detailed description of the task and scoring rubric is also provided.
+In addition to human evaluation, we also use an LLM (GPT-4-32K) for evaluating the battles as well as providing scores using the direct assessment metrics.
 3.3.1 Pairwise comparison
@@
 The annotators are provided with the query, the two model responses (model names are hidden), and set of three options - A (response 1 is better), B (response 2 is better), and C (tie, equally good/bad).
-We also ask the annotators to provide a spoken justification for the chosen response that is captured as audio by the app.
+During evaluation, we ask the annotators to provide a spoken justification for the chosen response that is captured as audio by the app.
 The annotation guidelines and Hindi app screenshots are available in Appendix E.1.
 LLM Evaluation Setup
-We also evaluate battles using GPT-4-32K as an LLM evaluator. The setting is similar to the one provided to humans. The detailed prompt is provided in Figure 10.
+We also evaluate battles using GPT-4-32K as an LLM evaluator.
+The setting is similar to the one provided to humans with the same rubric and a similar set of instructions provided as a prompt to the LLM.
+The detailed prompt is provided in Figure 10.
 3.3.2 Direct Assessment
@@
 3.3.3 Safety Evaluation
-We use the Hindi prompts from RTP-LX de Wynter et al.
-(2024) dataset [^3] https://github.com/microsoft/RTP-LX which is specifically designed to elicit toxic responses and ask the relevant models to generate completions.
-These completions are then evaluated using an LLM evaluator with the same prompt used for individual evaluations (Fig 11).
+We also conduct a preliminary safety evaluation to estimate the efficacy of guardrails in different LLMs.
+For this, we use the Hindi prompts from RTP-LX [^4] https://github.com/microsoft/RTP-LX de Wynter et al.
+(2024) which is specifically designed to elicit toxic responses and ask the models under consideration to generate completions.
+These completions are then evaluated using an LLM evaluator with the same prompt used for individual evaluations (Figure 11).
+We do not employ humans for the safety evaluation due to ethical concerns.
 The detailed rubric for Safety is defined in Fig 15.
-We also perform an exact match with the Hindi block words from the FLORES Toxicity-200 dataset [^4] https://github.com/facebookresearch/flores/blob/main/toxicity/README.md Costa-jussà et al.
+We also perform an exact match with the Hindi block words from the FLORES Toxicity-200555https://github.com/facebookresearch/flores/blob/main/toxicity/README.md Costa-jussà et al.
 (2022) to check for toxic words in the output.
@@
 To check for the quality of human annotation, we calculate inter-annotator agreement between the three human annotators using two metrics - Percentage Agreement (PA) and Fleiss Kappa ($\kappa$).
-These metrics are also used to judge the alignment between humans and LLMs for the evaluation tasks, following the same setup as prior work Hada et al.
+These metrics are also used to judge the alignment between humans and LLMs for the evaluation tasks, following the same setup as our prior work Hada et al.
 (2024b, a).
@@
 For the safety evaluation, the scores depict the fraction of prompts for which models gave problematic content.
-A detailed description of how each leaderboard is constructed along with the scores is available in Appendix G.
+A detailed description of how each leaderboard is constructed along with the scores is available in Appendix F.
 Figure 2: Comparison of Elo ratings of models across languages evaluated by both humans and an LLM.
@@
 An interesting next step would be to evaluate fine-tuned versions of Llama-3 70B, as and when they are available.
-Proprietary LLMs like GPT-4 and Gemini-Pro 1.0555available only for Hindi and Bengali rank in the upper middle portion of the human evaluations leaderboard, and top most of the LLM-evaluator leaderboards, showing evidence of self-bias by GPT-4 Panickssery et al.
+Proprietary LLMs like GPT-4 and Gemini-Pro 1.0666available only for Hindi and Bengali rank in the upper middle portion of the human evaluations leaderboard, and top most of the LLM-evaluator leaderboards, showing evidence of self-bias by GPT-4 Panickssery et al.
 (2024); Xu et al.
 (2024).
-GPT-3.5-Turbo, however, ranks across the lower-middle half and performs worse than the finetuned Indic LLMs.
+GPT-3.5-Turbo, however, ranks across the lower-middle half and performs worse than the fine-tuned Indic LLMs.
 The LLM evaluator also tends to favour Gemma 7B more than humans, suggesting that there may be some artifacts in some models that the LLM-evaluator picks up on.
@@
 4.2 RTP-LX Safety Analysis
-We conduct a safety analysis of all the Hindi LLMs, and evaluate the completions using GPT-4-32K.
+We present the preliminary safety analysis of all the Hindi LLMs, for which we evaluate the completions using GPT-4-32K.
 Following Hada et al.
-(2024a), we use a temperature of 1.0 to elicit as lower-probability generations as possible, which might be problematic.
+(2024a), we use a temperature of 1.0 to elicit even unexpected generations which might be problematic.
 API based LLMs, such as GPT and Gemini-Pro usually have guardrails and content moderation services before the actual model, and hence, we find that our prompts are blocked.
@@
 GPT-3.5-Turbo produces the least toxic completions ($\sim$ 10%), followed by GPT-4o and GPT-4.
-The AryaBhatta-Gemma models produce the highest number of toxic completions ($\sim$ 60%), while the Aya Model generated the most toxic/profane content in its generations upon manual checking.
+We also note that the Gemma model and its fine tuned variants (Aryabhatta variants and Navarasa) consistently perform the worst.
+However, since these evaluations are automated, they may contain potential biases.
+We leave further study with human evaluations as part of future work.
 Figure 4: RTP-LX Safety Evaluation of Hindi models. We report the fraction of prompt completions judged problematic by GPT-4 Evaluator and the heuristic Toxicity-200 exact match.
@@
 We compute the Percentage Agreement (PA) and Fleiss Kappa ($\kappa$) score which are calculated at a per-datapoint level as well as the general agreement between the leaderboards using Kendall’s Tau ($\tau$).
-The PA score is reported in Appendix H.
-Query Type Pairwise Direct
+The PA score is reported in Appendix G.
+Prompt Type Pairwise Direct
 […]
@@
 Non-Cultural 0.57 0.55 0.49 0.37
-Table 1: Average Fleiss Kappa ($\kappa$) correlations between Humans and Human-LLM for both evaluations across query types. Here $\mathcal{H}$ stands for Humans.
+Table 3: Average Fleiss Kappa ($\kappa$) correlations between Humans and Human-LLM for both evaluations across prompt types. Here $\mathcal{H}$ stands for Humans.
 Pairwise Battles
 On average humans have a moderate $\kappa$ score [^8] Generally, $\kappa>0.45$ is considered strong positive agreement of 0.54 whereas the human-average and LLM have a $\kappa$ score of 0.49.
-An ablation on the query-type in Table 1 reveals that LLM evaluator agrees comparatively less on the culturally-nuanced queries.
+An ablation on the prompt-type in Table 3 reveals that LLM evaluator agrees comparatively less on the culturally-nuanced prompt.
 A language-wise breakdown of the $\kappa$ scores can be seen in Figure 5.
@@
 In this case, humans tend to have a slightly lower but similar agreement to the pairwise scores.
-However, the agreement between humans and LLMs significantly drops and is the lowest again for the culturally-nuanced set of queries.
+However, the agreement between humans and LLMs significantly drops and is the lowest again for the culturally-nuanced set of prompts.
 Figure 5 shows that for Direct Assessment, humans have similar agreement across all languages but the human-LLM agreement is significantly lower particularly for Bengali and Odia.
@@
 We check the agreement between the leaderboards to get a sense of agreement on higher level trends.
-We report the Kendall Tau ($\tau$) scores between the rankings of models in both Elo and Direct Assessment leaderboards in Table 36.
+We report the Kendall Tau ($\tau$) scores between the rankings of models in both Pairwise (Elo) and Direct Assessment (DA) leaderboards in Table 4.
 On average we see a high $\tau$ score [^9] Generally, $\tau>0.7$ is considered a strong positive correlation of 0.76 for the Elo leaderboards which signifies that the human and LLM-evaluator agree on the general trends.
-From Figure 2 we can see that although the absolute rankings are not same, we can still find similar sub-group of models.
-We see this agreement go down for the Direct Assessment leaderboard which has an average agreement of 0.65.
-This reinforces our hypothesis that the LLM evaluator is worse at the Direct Assessment task.
+From Figure 2 also we can see that although the absolute rankings are not same, we can still find similar sub-group of models.
+We see this agreement go down for the DA leaderboard which has an average agreement of 0.65.
+We also report a detailed language-wise breakdown and find that for Elo leaderboards, Gujarati has the highest correlation while Bengali is the lowest.
+For the DA leaderboards, the scores drop and it is again the lowest for Bengali.
+This reinforces our hypothesis that the LLM evaluator is worse at the DA task.
+Language Pairwise Direct Average 0.76 0.65 Bengali 0.66 0.43 Gujarati 0.85 0.75 Hindi 0.80 0.67 Kannada 0.76 0.55 Malayalam 0.82 0.66 Marathi 0.82 0.82 Odia 0.78 0.53 Punjabi 0.69 0.54 Tamil 0.71 0.60 Telugu 0.70 0.91
+Table 4: Kendall Tau ($\tau$) correlations between Pairwise (Elo) and Direct Assessment leaderboards constructed through human annotators and LLM evaluator.
 4.4 Bias Analysis
@@
 We also analyse if there is any bias by humans or LLM evaluator to pick a longer response as the better one.
-First, we show the distribution of the average length of winning, losing and tied responses for pairwise evaluation in Figure 9.
-We can see that both the winning and losing responses have similar distributions with a median length of approximately 80 words for both evaluators.
-Second, we also investigate the correlation between response length and the Direct Assessment scores and again find no such correlations for both the evaluators.
-Hence, we find no significant evidence of bias towards the length of responses against the findings of Wu and Aji (2023).
-Figure 9: Response length distributions for the winning, losing and tied responses by humans and LLM evaluator in pairwise evaluations.
+From Figure 9, we observe that both humans and LLMs exhibit a slight bias towards a longer response which increases with the size difference between the responses.
+For LLMs, the bias remains relatively stable when the difference is 40-100 words, hovering around 0.6.
+Additionally, we observe that humans are slightly more biased than LLMs.
+We also note that the bias decreases when the difference becomes too large (over 100 words).
+We limit our responses to 300 words and hypothesize that such large responses might contain gibberish since LLMs are not very adept in multilingual settings.
+Hence, we find evidences of bias towards the length of responses by both the evaluators Wu and Aji (2023).
+Figure 9: Figure showing the win fraction of a longer answer over a shorter answer (0.5 indicates random) against the difference in length of the responses in pairwise comparisons.
 4.4.4 Self-Bias
 Lastly, we check for self-bias by the GPT evaluator towards its own outputs across both types of evaluation.
-We calculate the average rank of GPT-4 in the Elo leaderboard given by both the evaluators as well as the rank of all other models which were evaluated for all the 10 languages.
-We find that of the 10 selected models, the average rank of GPT-4 increases by the highest amount (1.4 places) for evaluations performed by the GPT evaluator.
+We calculate the average rank of GPT-4 in the Elo leaderboard given by both the evaluators as well as the rank of all other models which were evaluated for atleast 8-10 languages.
+We find that of the 11 selected models, the average rank of GPT-4 increases by the highest amount (1.4 places) for evaluations performed by the GPT evaluator (Appendix H).
 The rationale behind using the Elo leaderboard is that simply checking the win-rate would not account for biases due to GPT evaluator giving lesser ties than humans.
@@
 This section summarizes the feedback from three annotators per language on their experiences with pairwise comparison and direct assessment tasks.
-Q1 Were the annotators able to understand the pairwise evaluation task and what problems did they face?
+Q1. Were the annotators able to understand the pairwise evaluation task and what problems did they face?
 Majority of the annotators were able to understand the guidelines clearly and found the task simple.
-They also noted that linguistic acceptability played a big role in determining the easiness of the task, languages like Odia and Tamil had many grammatical errors which made it a difficult to go through responses.
-Q2 Were the annotators able to understand the metrics in direct assessment?
+They also noted that linguistic acceptability played a big role in determining the easiness of the task, languages like Odia and Tamil had many grammatical errors which made it difficult to go through responses.
+Q2. Were the annotators able to understand the metrics in direct assessment?
 The annotators found this task moderately difficult and some of them found the hallucination concept a bit tricky to understand.
 This task required the annotators to do online-search to check for hallucinations which made it more time-consuming.
-Q3 Which type of evaluation did the annotators find easier?
+Q3. Which type of evaluation did the annotators find easier?
 Most annotators found the pairwise comparison task easier in comparison to the direct assessment since it did not require them to evaluate every aspect of a response in detail and was less time-consuming.
@@
 5 Discussion
-Multilingual Performance
+RQ1. Are Indic LLMs able to compete with Proprietary models in respective languages?
 From our evaluations, we find that smaller Indic models perform better than the open-source models they are trained on, and larger frontier models such as GPT-4o perform best on Indic languages.
@@
 Our evaluation not only provides a ranking of LLMs but also indicates which open source models (like Llama-3) are potentially promising starting points for fine-tuning language specific Indic models.
-Human-LLM Agreement
-We find that LLM evaluators agree fairly well with the humans on the pairwise evaluation task in comparison to the direct assessment task.
+RQ2. Can LLM evaluators be used as a substitute for human evaluators in the multilingual setting? Which task format shows higher promise?
+We find that LLM evaluators agree fairly well with humans on the pairwise evaluation task in comparison to the direct assessment task.
 The LLM evaluator has low agreement with humans on Marathi, Bengali and Punjabi in the pairwise task and very low agreement for all languages particularly Bengali and Odia in direct assessment task.
@@
 On manually going through some examples, we find that humans tend to prefer outputs which are more friendly in nature, i.e., have good formatting, use colloquial language and explain with the help of examples.
-We find that LLM-evaluators agree less with humans on evaluating responses with cultural nuances, suggesting that they do no possess enough cultural context to do these kinds of evaluations well.
+We find that LLM-evaluators agree less with humans on evaluating responses with cultural-nuances, suggesting that they do no possess enough cultural context to do these kinds of evaluations well.
 However, LLM evaluators are still able to capture general trends at a higher level as seen from the $\tau$ scores.
 This suggests that a human-in-the-loop or hybrid evaluation system is necessary for performing multilingual, multi-cultural evaluation.
-Biases in Judgement
+RQ3. What are the biases that affects the evaluators’ judgements?
 We look for position bias by looking at evaluator behaviour on option flipping and find no such biases.
@@
 This leads to LLM evaluators having higher scores in the direct assessment task as well as fewer ties (more decisive) in pairwise evaluations task.
-We also look for correlations between response length and a winning response and again find no such bias.
+We also look for correlations between response length and a winning response and find slight bias towards a longer response by both evaluators.
 Lastly, we check for any self-bias in the GPT evaluator and find evidence of it preferring its own output.
@@
 Appendix B Model Details
-We list the details of all the models evaluated by us in our study in Table 2 and Table 3. We conduct evaluations on 20 indic and 10 multilingual models.
+We list the details of all the models evaluated by us in our study in Table 5 and Table 6.
+We conduct evaluations on 20 indic and 10 multilingual models.
+The models are classified based on the following criteria,
+  1. 1.
+    Indic: Model fine-tuned specifically for only Indian language(s).
+  2. 2.
+    Open-Source: Openly available models not fine-tuned specifically for Indian languages.
+  3. 3.
+    Proprietary: Close source models with API access and no information about the model/training data.
 Model Short Name Model Type Hindi Models ai4bharat/Airavata Gala et al.
@@
 All evaluated models are prompted with a system instruction followed by the query with no few-shot examples.
-The prompt template for each open-source model is taken from their HuggingFace model card wherever applicable, else the default Llama2-prompt is used.
+The prompt template for each open-source model is taken from their HuggingFace model card wherever applicable, else the default Llama2-prompt i.e., [INST]<SYS>, is used.
 We instruct the models to limit their responses to 300 words and truncate the responses when necessary to make human evaluation easier, as Karya  workers perform the evaluation tasks on a smartphone.
@@
 {promptbox}
-‘‘name": ‘‘task quality",
+‘‘name": ‘‘task_quality",
 ‘‘description": ‘‘Task Quality gauges the degree to which a model adheres to and executes the specific directives given in the prompt.
@@
 D.3 RTP-LX Safety Evaluation
-We also do a safety evaluation study for the Hindi models using the RTP-LX de Wynter et al.
+We also conduct a preliminary safety evaluation study for the Hindi models using the RTP-LX de Wynter et al.
 (2024) Hindi dataset.
 Only LLM evaluators were used for this study.
-We used the same instruction prompt as used in direct assessment above (Figure 11 and calculate the problematic content score in the model output generations.
+We used the same instruction prompt as used in direct assessment above (Figure 11) and calculate the problematic content score in the model output generations.
 The problematic content rubric can be seen in Figure 15.
@@
 Figure 16: Detailed task instructions provided to the annotators.
-Figure 17: Hindi App screenshots for pairwise human evaluations
+Figure 17: Hindi App screenshots for pairwise human evaluations.
 E.2 Direct Assessment
@@
 Figure 19: Hindi App screenshots for Direct Assessment human evaluations.
-Appendix F Evaluation Statistics
-We perform a total of over 90k human evaluations and 30k LLM evaluations in our study.
-A breakdown of these statistics can be seen in Table 4.
-A total of 21690 datapoints (battles) are evaluated by LLMs evaluator for pairwise evaluation whereas 2880 model query-response pairs are evaluated for 3 metrics (hallucinations, task quality and linguistic acceptability) which results to 8640 datapoints in total.
-Hence, a total of $21690+8640=30330$ evaluations are performed by the LLM evaluator and since we annotate each datapoint by 3 humans, a total of 90990 evaluations are done by humans.
-In the models column, first number within parenthesis is the number of Indic-only models and the second value is the number of multilingual models under evaluation.
-We evaluate 12-15 models for each language (except Hindi) and 20 models for Hindi.
-Language Models Pairwise Direct All 30 (20+10) 21690 8640 Hindi 20 (10+10) 4180 1200 Telugu 15 (7+8) 2310 900 Bengali 15 (6+9) 2310 900 Malayalam 14 (6+8) 2002 840 Kannada 14 (6+8) 2002 840 Tamil 14 (6+8) 2002 840 Odia 14 (6+8) 2002 840 Gujarati 13 (5+8) 1715 780 Punjabi 13 (5+8) 1715 780 Marathi 12 (4+8) 1452 720
-Table 4: Number of pairwise comparison (battle) and direct assessment datapoints for each language.
-Both LLM evaluator and Humans were used for all datapoints.
-In the models column, first number within parenthesis is the number of Indic-only models and the second value is the number of multilingual models under evaluation.
-Total evaluations: $21690+8640=30330$ for LLM, and $3\times 30330=90990$ for humans, as each data point was annotated by 3 humans.
-Appendix G Leaderboards
+Appendix F Leaderboards
 In this section we present the detailed leaderboards constructed by the strategies discussed in Section 3.
@@
 This is done due to the lower number of datapoints and to get confidence intervals.
-G.1 MLE Elo Leaderboards
-We report the MLE Elo leaderboards for all the 10 languages in Tables 5, 6, 7, 8, 9, 10, 11, 12, 13 and 14.
+F.1 MLE Elo Leaderboards
+We report the MLE Elo leaderboards for all the 10 languages in Tables 7, 8, 9, 10, 11, 12, 13, 14, 15 and 16.
 Model Rank (Human) Elo Rating (Human) Rank (LLM) Elo Rating (LLM) GPT-4o 1 1551 $\pm$ 18.95 3 1604 $\pm$ 22.73 Llama-3 70B 2 1444 $\pm$ 13.09 5 1538 $\pm$ 18.09 Gemini-Pro 1.0 3 1444 $\pm$ 15.93 2 1672 $\pm$ 21.87 GPT-4 4 1346 $\pm$ 12.59 4 1598 $\pm$ 20.44 SamwaadLLM 5 1247 $\pm$ 11.98 1 1688 $\pm$ 21.51 Llama-3 8B 6 1116 $\pm$ 12.37 6 1233 $\pm$ 16.0 Navarasa 7 1095 $\pm$ 12.27 11 955 $\pm$ 12.85 AryaBhatta-GemmaOrca 8 1067 $\pm$ 10.7 10 975 $\pm$ 12.91 AryaBhatta-Llama3GenZ 9 1066 $\pm$ 10.17 7 1157 $\pm$ 14.33 GPT-3.5-Turbo 10 1053 $\pm$ 10.71 8 1086 $\pm$ 13.49 AryaBhatta-GemmaUltra 11 1025 $\pm$ 10.88 12 935 $\pm$ 13.08 OdiaGenAI-Bengali 12 860 $\pm$ 9.39 15 719 $\pm$ 11.09 Gemma 7B 13 859 $\pm$ 9.29 9 1029 $\pm$ 14.42 Mistral 7B 14 821 $\pm$ 8.97 13 891 $\pm$ 12.85 Llama-2 7B 15 800 $\pm$ 0.0 14 800 $\pm$ 0.0
@@
 Table 16: MLE Elo for Telugu
-G.2 Standard Elo Leaderboards
-We report the Standard Elo leaderboards for all the 10 languages in Tables 15, 16, 17, 18, 19, 20, 21, 22, 23 and 24.
+F.2 Standard Elo Leaderboards
+We report the Standard Elo leaderboards for all the 10 languages in Tables 17, 18, 19, 20, 21, 22, 23, 24, 25 and 26.
 Model Rank (Human) Elo Rating (Human) Rank (LLM) Elo Rating (LLM) GPT-4o 1 1522 $\pm$ 21.79 3 1499 $\pm$ 22.77 Llama-3 70B 2 1422 $\pm$ 17.19 5 1444 $\pm$ 19.07 Gemini-Pro 1.0 3 1420 $\pm$ 21.5 2 1557 $\pm$ 18.81 GPT-4 4 1328 $\pm$ 19.92 4 1493 $\pm$ 23.95 SamwaadLLM 5 1233 $\pm$ 17.95 1 1569 $\pm$ 22.26 Llama-3 8B 6 1107 $\pm$ 18.05 6 1194 $\pm$ 19.26 Navarasa 7 1088 $\pm$ 20.58 11 939 $\pm$ 19.32 AryaBhatta-GemmaOrca 8 1061 $\pm$ 18.7 10 957 $\pm$ 18.27 AryaBhatta-Llama3GenZ 9 1057 $\pm$ 18.08 7 1126 $\pm$ 22.24 GPT-3.5-Turbo 10 1046 $\pm$ 17.03 8 1065 $\pm$ 21.33 AryaBhatta-GemmaUltra 11 1020 $\pm$ 17.28 12 920 $\pm$ 21.15 Gemma 7B 12 860 $\pm$ 14.8 9 1006 $\pm$ 20.86 OdiaGenAI-Bengali 13 859 $\pm$ 16.36 15 723 $\pm$ 14.98 Mistral 7B 14 820 $\pm$ 13.87 13 881 $\pm$ 18.95 Llama-2 7B 15 800 $\pm$ 0.0 14 800 $\pm$ 0.0
@@
 Table 26: Standard Elo for Telugu
-G.3 Direct Assessment Leaderboards
-We report the Direct Assessment leaderboards for all the 10 languages in Tables 25, 26, 27, 28, 29, 30, 31, 32, 33 and 34.
+F.3 Direct Assessment Leaderboards
+We report the Direct Assessment leaderboards for all the 10 languages in Tables 27, 28, 29, 30, 31, 32, 33, 34, 35 and 36.
 Model Rank (Human) LA (Human) TQ (Human) H (Human) Score (Human) Rank (LLM) LA (LLM) TQ (LLM) H (LLM) Score (LLM) GPT-4o 1 1.80 1.75 0.80 4.35 1 2 2 1 5 Llama-3 70B 2 1.70 1.60 0.75 4.05 1 2 2 1 5 Gemini-Pro 1.0 3 1.35 1.40 0.60 3.35 1 2 2 1 5 AryaBhatta-GemmaOrca 4 1.15 0.85 0.60 2.60 10 1.20 1.35 0.65 3.20 Navarasa 5 1.05 0.90 0.55 2.50 12 1 1.15 0.65 2.80 AryaBhatta-GemmaUltra 6 1.15 0.80 0.50 2.45 11 1.10 1.15 0.65 2.90 GPT-3.5-Turbo 7 1.05 0.90 0.30 2.25 7 1.90 1.95 0.95 4.80 Llama-3 8B 8 1.25 0.70 0.30 2.25 8 1.85 1.85 0.95 4.65 GPT-4 9 0.80 0.95 0.40 2.15 1 2 2 1 5 SamwaadLLM 10 0.85 0.80 0.40 2.05 1 2 2 1 5 AryaBhatta-Llama3GenZ 11 0.80 0.70 0.40 1.90 6 2 1.95 1 4.95 Gemma 7B 12 0.25 0.10 0.05 0.40 9 1.70 1.75 0.80 4.25 OdiaGenAI-Bengali 13 0.25 0.05 0.05 0.35 14 0.25 0.20 0.05 0.50 Mistral 7B 14 0.05 0.05 0 0.10 13 1.25 1.15 0.35 2.75 Llama-2 7B 15 0 0 0 0 15 0.10 0.05 0 0.15
@@
 Table 36: Direct Assessment Leaderboard for Telugu
-Appendix H Detailed Agreement Scores
-H.1 Percentage Agreement
+Appendix G Percentage Agreement
 In this section we report the Percentage Agreement (PA) scores which gives a raw-interpretable number but does not take class-imbalance into account. The scores are reported in Table 37.
 Pairwise Battles
-On average humans agree on 70% of the samples among themselves whereas the accuracy is similar albeit slightly lower for humans and LLM evaluator.
+On average humans agree on 70% of the samples among themselves whereas the accuracy is similar albeit slightly lower for human-average and LLM evaluator.
 We see both evaluators agree more on the non-cultural subset of queries and follow a similar trend to the Fleiss Kappa correlations reported in Table 3.
@@
 Direct Assessment
-For this task, we find slightly higher agreement between humans in comparison to the pairwise evaluation and it is similar for both the query_types.
+For this task, we find slightly higher agreement between humans in comparison to the pairwise evaluation and it is similar for both the prompt types.
 However we see a decline between human and LLM evaluator agreement and it is the worse for culturally-nuanced set of queries.
@@
 Figure 20: Language-wise PA scores breakdown for Pairwise and Direct Assessment evaluations.
-H.2 Kendall Tau
-We also report a detailed language-wise breakdown of the Kendall Tau rank correlations for both evaluation tasks in Table 36.
-We find the correlation between human and LLM leaderboards to be higher for the pairwise evaluation.
-Gujarati has the highest correlation while Bengali is the lowest.
-For the direct assessment the scores drops and it is again the lowest for Bengali.
-We conclude that evaluators are better at capturing general trends in a pairwise comparison evaluation over a direct assessment task.
-Language Pairwise Direct Average 0.76 0.65 Bengali 0.66 0.43 Gujarati 0.85 0.75 Hindi 0.80 0.67 Kannada 0.76 0.55 Malayalam 0.82 0.66 Marathi 0.82 0.82 Odia 0.78 0.53 Punjabi 0.69 0.54 Tamil 0.71 0.60 Telugu 0.70 0.91
-Table 36: Kendall Tau ($\tau$) correlations between Pairwise (Elo) and Direct Assessment leaderboards constructed through human annotators and LLM evaluator.
+Appendix H Self-Bias
+In this section we present detailed results of the self-bias analysis.
+Table 38 shows the average change in Elo leaderboard ranks across languages when evaluated by GPT-4 evaluator in comparison to humans.
+We select the 11 models which are evaluated on atleast 8-10 out of the total 10 languages.
+We see GPT variants rank increase the most.
+Gemma also shows an increase while the Aryabhatta variants and Llama-3 70B model have a big drop.
+S. No. Model $\Delta$ Rank
+1 GPT-4 +1.4
+2 Gemma 7B +1.3
+3 GPT-4o +0.6
+4 GPT-35-Turbo +0.4
+5 Llama-3 8B +0.1
+5 Llama-2 7B +0.1
+7 Mistral 7B -0.4
+8 Navarasa -0.5
+9 Arybhatta-GemmaOrca -1.3
+10 Llama-3 70B -1.6
+11 Arybhatta-GemmaUltra -1.9
+Table 38: Average change in Elo Rank ($\Delta$) across languages when evaluated by GPT-evaluator in comparison to humans.